### v1

# Enumeration of maps with tight boundaries and the Zhukovsky transformation

Abstract
We consider maps with tight boundaries, i.e. maps whose boundaries have minimal length in their homotopy class, and discuss the properties of their generating functions $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ for fixed genus $g$ and prescribed boundary lengths $\ell_{1},\ldots,\ell_{n}$, with a control on the degrees of inner faces. We find that these series appear as coefficients in the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$, a fundamental quantity in the Eynard-Orantin theory of topological recursion, thereby providing a combinatorial interpretation of the Zhukovsky transformation used in this context. This interpretation results from the so-called trumpet decomposition of maps with arbitrary boundaries. In the planar bipartite case, we obtain a fully explicit formula for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ from the Collet-Fusy formula. We also find recursion relations satisfied by $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, which consist in adding an extra tight boundary, keeping the genus $g$ fixed. Building on a result of Norbury and Scott, we show that $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ is equal to a parity-dependent quasi-polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ times a simple power of the basic generating function $R$. In passing, we provide a bijective derivation in the case $(g,n)=(0,3)$, generalizing a recent construction of ours to the non bipartite case.
1 Introduction
1.1 Context and motivations
We pursue our investigation, started in [triskell, polytightmaps], of the enumerative properties of maps with tight boundaries. In this episode, we add a new twist to the story, by making an unexpected connection with the theory of topological recursion [Eynard2007, Eynard2016]. In particular, we find a combinatorial interpretation of the Zhukovsky transformation, which was so far used as an analytical tool in this theory.
Let us first recall some context. The enumeration of maps (graphs embedded in surfaces) is a venerable topic in combinatorics, initiated by Tutte in his famous series of “Census” papers, see [Tutte1963] and references therein. Accounts of further developments may be found for instance in [Goulden2004, Schaeffer15]. A fertile connection with random matrices was made in [BIPZ], and it is essentially its ramifications, motivated by the study of 2D quantum gravity—see e.g. [Witten1991, DiFrancesco1995, Ambjoern1997] for overviews—that led to topological recursion.
In the paper [triskell], we found a surprisingly simple formula for the generating function of planar bipartite maps with three tight boundaries (as we will explain in more detail below, a boundary is said tight if it has minimal length in its homotopy class). The tightness property seems to play a role, as without it the corresponding generating function becomes slightly more involved [CoFu12]. One may therefore wonder whether a similar phenomenon occurs for maps of other topologies (higher genus, more boundaries).
This question was already explored in the limit case of tight maps. Colloquially speaking, a tight map is a map in which every face is seen as a boundary, and is forced to be tight. The counting of tight maps was actually investigated first by Norbury [Norbury2010], as tight maps are nothing but fatgraphs describing lattice points in the moduli space of curves. A remarkable quasi-polynomiality phenomenon occurs in this problem: if we denote by $N_{g,n}(b_{1},\ldots,b_{n})$ the number of tight maps of genus $g$ with $n$ boundaries of lengths $b_{1},\ldots,b_{n}$, then Norbury showed that it is a quasi-polynomial of degree $3g-3+n$ in $b_{1}^{2},\ldots,b_{n}^{2}$ depending on the parities, and we gave in [polytightmaps] a bijective construction of these quasi-polynomials in the planar case $g=0$.
In a further paper [NoSc13], Norbury and Scott observed that the quasi-polynomiality phenomenon can be related to a specific feature in topological recursion, namely that the “$x$” meromorphic function attached to the “spectral curve” takes the particular form
$x(z):=\alpha+\gamma\left(z+z^{-1}\right).$ (1)
This form is nothing but the so-called Zhukovsky transformation [Zhukovskii1910]. The case of tight maps corresponds to taking the second “$y$” function equal to the Zhukovsky variable $z$, but Norbury and Scott showed that the quasi-polynomiality subsists when $y$ is modified.
In this paper, we relate the observation of Norbury and Scott to the enumeration of maps with tight boundaries. At the combinatorial level, the Zhukovsky transformation $z\mapsto x(z)$ corresponds to a substitution: namely we find that it translates the natural idea of transforming a map with tight boundaries into a map with arbitrary boundaries by gluing a “trumpet” onto each boundary (see Figure 2 below for an illustration). By trumpet, we mean a planar map with two boundaries, one of them being tight while the other is arbitrary. We use the term by analogy with [Saad2019], where a similar notion is introduced in the context of hyperbolic geometry. The coefficients $\alpha$ and $\gamma$ appearing in (1) are related to the “basic” generating functions $R$ and $S$ defined in the next section by $\alpha=S$ and $\gamma=R^{1/2}$.
By this approach, we deduce the interesting property that the generating function $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ of maps of genus $g$ and $n$ tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$ has the general form
$T^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}\gamma^{\ell_{1}+\cdots+\ell_{n}}$ (2)
with $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ a quasi-polynomial of degree $3g-3+n$ in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ depending on the parities, and $\gamma=R^{1/2}$ as before. In particular, in the planar bipartite case, we will derive a fully explicit expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ from the Collet-Fusy formula [CoFu12].
Let us conclude this section by mentioning the recent paper [Budd2023] which introduces the notion of tight boundaries in the context of hyperbolic geometry, and makes the connection with the so-called JT gravity [Saad2019]. A very similar polynomiality phenomenon occurs there, which makes us believe that we are looking at a same reality from different angles.
1.2 Basic definitions
Let us now introduce precisely the main definitions and conventions used in this paper.
Combinatorial notions: maps with boundaries, and their automorphisms.
For $g$ a nonnegative integer, a map of genus $g$ is a cellular embedding of a finite connected multigraph into the closed orientable surface of genus $g$, considered up to isomorphism. By cellular embedding, we mean that the graph (which consists of vertices and edges) is drawn on the surface without edge crossings, and that the connected components of the complement of the graph, called faces, are homeomorphic to open disks. A map of genus zero is said planar.
A map with boundaries is a map in which we mark some faces and vertices and call them boundaries. We shall use the terms boundary-face and boundary-vertex when we want to specify the type of a boundary. A face which is not a boundary is called an inner face. The length of a boundary is by definition equal to its degree (number of incident edge sides) for a boundary-face, and to 00 for a boundary-vertex. A boundary-face is said rooted if one of its incident corners is marked. A boundary-vertex cannot be rooted.
An automorphism of a map with boundaries is a graph isomorphism that also preserves the map structure, in the following sense.
  • The images of the edges pointing away from a given vertex, and listed cyclically in clockwise order, are the edges pointing away from the image vertex, also in cyclic clockwise order. In particular, the graph isomorphism also induces a bijection of the set of faces onto itself, respecting the incidence relations with vertices and edges.
  • The marked elements are preserved.
The automorphisms of a given map with boundaries form a group, which is often reduced to the identity element. For instance, this holds for a rooted map of any genus, or for a planar map possessing at least three distinguished boundaries. Contrary to the former situation, the latter is not true in higher genera: Figure 1 provides a counterexample in genus $1$.
Figure 1: A toric map, with three distinguished boundary-faces of degree 6, admitting an automorphism group of order $3$. There are only $2$ rooted maps resulting from marking one of the $6$ corners incident to the cyan face, and we must account for this fact by weighing this map by the inverse automorphism group order factor $1/3$.
Topological notions: boundaries and homotopy.
We think of boundaries as representing holes in the surface. Precisely, we create a puncture inside each boundary-face, while we draw a small circle around each boundary-vertex: the interior of the circle is removed but the edges incident to the vertex remain connected by the circle. See [triskell, Figure 2] for an illustration.
Given a map, possibly with boundaries, a path is a sequence of consecutive edges. It defines a curve on the underlying surface provided that, if the path visits a boundary-vertex, we choose a direction for “circumventing” the corresponding puncture in the surface. A path will always be assumed to contain this data in the following. A path is said closed if it starts and ends at the same vertex. A path is said simple if it does not visit the same vertex twice (except at its endpoints for a simple closed path). The contour of a face is the closed path formed by its incident edges. Two closed paths on the map are said (freely) homotopic to one another if their corresponding curves can be continuously deformed into one another on the punctured surface. A closed path is said homotopic to a boundary if, on the underlying surface of the map, the curve associated with the path can be contracted onto the puncture associated with the boundary. A boundary-face is said tight (resp. strictly tight) if its degree is not larger (resp. is strictly smaller) than the number of edges of any other closed path homotopic to it. A boundary-vertex is considered as (strictly) tight by convention.
Generating functions.
Let $t,t_{1},t_{2},t_{3},\ldots$ be a collection of formal variables. We attach a weight $t$ to each vertex which is not a boundary, and a weight $t_{i}$ to each inner face of degree $i$, $i=1,2,3,\ldots$. The global weight of a map is the product of the weights of all its (non-boundary-) vertices and faces, multiplied by the inverse of the order of the automorphism group of the map. This last factor yields better combinatorial properties for the corresponding generating functions, see the caption of Figure 1 for a quick explanation.
We now introduce two important quantities denoted $R$ and $S$, which are formal power series in $t,t_{1},t_{2},t_{3},\ldots$. The first one $R$ is the generating function of planar maps with two distinguished boundaries of length $1$. The second one $S$ is that of planar maps with two boundaries, one of length $1$ and the other of length 00. It is known [^1] See for instance [census] for a proof in the dual language. Beware that $t$ is a weight per edge in this reference, one has to set $t_{i}=t\,g_{i}$ to match the notations. that, as formal power series, $R$ and $S$ are determined by the following equations:
$\begin{split}R&=t+\sum_{i\geq 1}t_{i}\ [z^{-1}]\left(z+S+\frac{R}{z}\right)^{i-1},\\ S&=\sum_{i\geq 1}t_{i}\ [z^{0}]\left(z+S+\frac{R}{z}\right)^{i-1}.\\ \end{split}$ (3)
A simplification occurs in the essentially bipartite case $t_{1}=t_{3}=t_{5}=\cdots=0$, which corresponds to imposing that every inner face has even degree. Indeed we then have $S=0$ and $R$ satisfies the single equation
$R=t+\sum_{j\geq 1}t_{2j}\binom{2j-1}{j}R^{j}.$ (4)
In this case, $R$ can also be understood as the generating function of bipartite planar maps with two unrooted boundaries, one of length $2$ and the other of length 00.
1.3 Outline
The remainder of this paper is organized as follows. Section 2 is devoted to the trumpet decomposition which relates maps with arbitrary boundaries to maps with tight boundaries. We first describe in Section 2.1 the bijection for maps with boundary-faces only, before analyzing its enumerative consequences in Section 2.2. The case of maps having also boundary-vertices is discussed in Section 2.3. In Section 2.4, we derive from the Collet-Fusy formula an explicit expression for the generating function of planar bipartite maps with tight boundaries. Section 3 is devoted to recursion relations satisfied by the generating functions of maps with tight boundaries. These recursion relations consist in adding an extra tight boundary, preserving the genus. We first discuss in Section 3.1 the addition of a boundary-vertex, and then that of a boundary-face in Section 3.2. In Section 3.3, we check that the explicit expression for planar bipartite maps found in Section 2.4 satisfies these recursion relations. Section 4 is devoted to the quasi-polynomiality phenomenon: we first consider maps with boundary-faces only in Section 4.1, revisiting the result of Norbury and Scott. We then treat the case of maps having also boundary-vertices in Section 4.2: the proof is by induction on the number of boundaries, using the recursion relations of Section 3 for the induction step and results from topological recursion for the initialization. For the sake of clarity, we begin with the case of maps with even face degrees. The extension to the general case follows the same strategy but requires extra technical steps which are offloaded to the appendices. Concluding remarks are gathered in Section 5. Additional material is contained in the appendices: Appendix A generalizes the formula of [triskell] for the generating function of planar maps with three tight boundaries (i.e.  “tight pair of pants”) to the non-bipartite case. Appendices B and C contain the proof of the induction step and the initialization step, respectively, for the quasi-polynomiality phenomenon. Appendix D gives a combinatorial expression for the derivatives of inverse functions of several variables, used in Appendix C.
2 The trumpet decomposition
In this section, we discuss a bijective decomposition of a map with boundaries into a map with tight boundaries and a collection of annular maps with one strictly tight boundary (a trumpet). This generalizes [triskell, Proposition 6.5], which dealt with planar maps with three boundaries, to arbitrary genera and number of boundaries.
2.1 The decomposition theorem
Let us introduce some notation. A trumpet is a map $M$ of genus 00 with two boundary-faces $f,F$, such that the boundary-face $F$ is rooted, and such that the boundary-face $f$, called the mouthpiece, is unrooted and strictly tight. Recall from Section 1.2 that this means that the contour of $f$ is the unique closed path of minimal length among all closed paths of $M$ that are (freely) homotopic to $f$. By considering a leftmost geodesic from the root corner of the face $F$ aimed towards the face $f$, we may canonically distinguish one corner incident to $f$. We state the following simple but crucial fact.
Lemma 2.1.
The contour of the boundary of the mouthpiece of a trumpet is simple.
This lemma is a consequence of the fact that the trumpet $M$ is drawn on a topological cylinder, and is characteristic of this topology: as can be seen for example on Figure 2, tight faces need not have simple boundaries in maps of arbitrary topology (precisely, of negative Euler characteristic $\chi=2-2g-n$, where $g$ is the genus and $n$ is the number of boundaries).
Proof.
We may assume that the trumpet $M$ is embedded in the plane, with the puncture of the mouthpiece $f$ set at point 00, and the puncture of the rooted face $F$ sent to infinity. We let $\mathbb{D}$ be the unit open disk in $\mathbb{R}^{2}$ and $\mathbb{S}^{1}=\partial\mathbb{D}$ be the unit circle.
The contour of $f$ is a closed path, which we orient arbitrarily and denote by $c$. Plainly, it admits a simple closed subpath $c^{\prime}$ not reduced to a single vertex, and we aim at showing that $c^{\prime}=c$. By the Jordan-Schoenflies theorem, $c^{\prime}$ cuts the plane into two domains $D$ (bounded) and $\mathbb{R}^{2}\setminus\overline{D}$ (unbounded), and moreover, $c^{\prime}$ viewed as a continuous injective mapping $\mathbb{S}^{1}\to\mathrm{Im}(c^{\prime})$ can be extended into a homeomorphism $h:\mathbb{R}^{2}\to\mathbb{R}^{2}$ sending $\mathbb{D}$ to $D$. If 00 does not belong to $D$, it follows $c^{\prime}$ is homotopic to the trivial path in the punctured plane $\mathbb{R}^{2}\setminus\{0\}$, so that $c$ is homotopic to a strictly shorter path, contradicting the tightness assumption of $f$. Therefore, it must be that $0\in D$, so that $c^{\prime}$ is homotopic to $\mathbb{S}^{1}$ and therefore generates the fundamental group $\pi_{1}(\mathbb{R}^{2}\setminus\{0\})\simeq\mathbb{Z}$. Since $c$, being the contour of the face $f$ that contains 00, is also a generator, this implies that $c^{\prime}$ is either homotopic to $c$ or to its reversed path. But the tightness of $f$ shows that the length of $c^{\prime}$ is at least that of $c$, and the only possibility is that $c^{\prime}$ is equal to $c$. ∎
Next, let $M_{0}$ be a map of genus $g$ with $n$ tight rooted boundary-faces $f^{0}_{1},\ldots,f^{0}_{n}$, and let $M_{i},1\leq i\leq n$ be a sequence of trumpets (a brassband), with respective boundary-faces $f_{i},F_{i}$, where $f_{i}$ denotes the mouthpiece. We assume that the length of $f_{i}$ and the length of $f_{i}^{0}$ are equal for $1\leq i\leq n$.
For every $i\in\{1,\ldots,n\}$, we orient the contour of the faces $f^{0}_{i},f_{i}$ in such a way that $f^{0}_{i}$ lies to the right, and $f_{i}$ lies to the left of their oriented contours. Then, since the boundary of $f_{i}$ is a simple closed path, and since $f_{i},f_{i}^{0}$ have same length, one may identify these boundaries, in such a way that the root of $f_{i}^{0}$ and the distinguished corner of $f_{i}$ are matched, by gluing edges sequentially as they appear in contour order, for our choice of orientation. This results in a map $M=\Phi(M_{0},M_{1},\ldots,M_{n})$ with genus $g$ and with the $n$ rooted boundary faces $F_{1},\ldots,F_{n}$ inherited from $M_{1},\ldots,M_{n}$.
Theorem 2.2 (Trumpet decomposition of maps).
Let $g,n$ be integers such that $g\geq 0$, $n\geq 1$, and $(g,n)\neq(0,1)$ or equivalently $\chi=2-2g-n\leq 0$, and let $L_{1},\ell_{1},\ldots,L_{n},\ell_{n}$ be fixed positive integers. Then, the mapping $\Phi$ is a bijection between
  • sequences $(M_{0},M_{1},\ldots,M_{n})$, where $M_{0}$ is a map of genus $g$ with $n$ rooted tight boundaries of lengths $\ell_{i},1\leq i\leq n$ and for $1\leq i\leq n$, $M_{i}$ is a trumpet with mouthpiece length $\ell_{i}$ and whose other external face has length $L_{i}$, and
  • maps $M$ of genus $g$ with $n$ rooted boundaries $F_{1},\ldots,F_{n}$ of lengths $L_{1},\ldots,L_{n}$, and such that for $1\leq i\leq n$, the minimal length of a closed path homotopic to $F_{i}$ is equal to $\ell_{i}$.
Figure 2: An illustration of the decomposition of a map of genus $3$ with three external faces into a map with tight boundaries and with the same topology, and a brassband of three trumpets. We emphasize on this picture that, while the mouthpieces of the trumpets always have simple boundaries by Lemma 2.1, the (tight) external faces of the central map need not have simple boundaries.
This result is illustrated in Figure 2. As already mentioned above, this is a generalization of [triskell, Proposition 6.5]. The main difficulty comes from constructing the reverse mapping $\Phi^{-1}(M)$, where $M$ is a map of genus $g$ with $n$ boundary-faces $F_{1},\ldots,F_{n}$. This consists in considering, for $i\in\{1,2,\ldots,n\}$, the set $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ of closed paths in $M$ that are homotopic to $F_{i}$, and that have minimal possible length, and to define an order $\prec^{(i)}$ that makes $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ a lattice. In words, we have $c\prec^{(i)}c^{\prime}$ if $c$ is closer to $F_{i}$ than $c^{\prime}$, although this description requires some interpretation, which is done in [triskell] by working in the universal cover of $M$. Since $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ is clearly a finite set, this implies that there is a smallest element $c^{i}$ in $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$, that we call the outermost minimal closed path homotopic to $F_{i}$. Cutting along these outermost closed paths, for every $i\in\{1,\ldots,n\}$, produces $n$ trumpets $M_{1},\ldots,M_{n}$ and a remaining map $M_{0}$ with $n$ tight boundaries $f^{0}_{1},\ldots,f^{0}_{n}$, whose lengths match with those of the corresponding mouthpieces.
We will not repeat the detailed argument here and refer the interested reader to [triskell, Section 6.1].
2.2 Enumerative consequences
Theorem 2.2 admits the following immediate corollary in terms of generating functions (recall that we attach a weight $t$ to each vertex which is not a boundary, and a weight $t_{i}$ to each inner face of degree $i$, $i=1,2,3,\ldots$).
Corollary 2.3.
For $g\geq 0$ and $L_{1},\ldots,L_{n}$ positive integers ($n\geq 1$), let us denote by $F^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of maps of genus $g$ with $n$ rooted boundary-faces of lengths $L_{1},\ldots,L_{n}$, and by $\hat{T}^{(g)}_{L_{1},\ldots,L_{n}}$ the generating function of those maps whose boundaries are tight. Then, for $(g,n)\neq(0,1)$ or equivalently $\chi=2-2g-n\leq 0$, we have
$F^{(g)}_{L_{1},\ldots,L_{n}}=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}\,,$ (5)
where $A_{L,\ell}$ is the generating function of trumpets with rooted boundary of length $L$ and mouthpiece of length $\ell$, where the vertices incident to the mouthpiece do not receive a weight $t$.
The usefulness of this statement comes from the fact that $A_{L,\ell}$ admits an explicit expression.
Proposition 2.4.
[irredmaps, Section 9.3] For any two positive integers $L,\ell$, we have
$A_{L,\ell}=[z^{\ell}]\left(z+S+\frac{R}{z}\right)^{L}\,.$ (6)
Note that the notation from [irredmaps] differs from that of the present paper: precisely, Proposition 2.4 is implied by formula (9.19) in this reference, in the case $d=0$, $d^{\prime}=\ell$ and $n=L$.
Let us now introduce the “grand” generating function
$W_{n}^{(g)}(x_{1},\ldots,x_{n}):=\sum_{L_{1},\ldots,L_{n}\geq 1}\frac{F^{(g)}_{L_{1},\ldots,L_{n}}}{x_{1}^{L_{1}+1}\cdots x_{n}^{L_{n}+1}}$ (7)
which is a formal power series in $x_{1}^{-1},\ldots,x_{n}^{-1}$ ($n\geq 1$). Then, following [Eynard2016, Definition 3.3.1, p. 87], we define for $(g,n)\neq(0,1),(0,2)$ [^2] Some conventional terms have to be added in these cases., or equivalently $\chi=2-2g-n<0$, the quantity
$\omega_{n}^{(g)}(z_{1},\ldots,z_{n}):=W_{n}^{(g)}\left(x(z_{1}),\ldots,x(z_{n})\right)x^{\prime}(z_{1})\cdots x^{\prime}(z_{n})$ (8)
where
$x(z):=R^{1/2}\left(z+z^{-1}\right)+S$ (9)
is precisely the Zhukovsky transformation discussed in Section 1.1.
Theorem 2.5 (Combinatorial interpretation of $\omega_{n}^{(g)}$).
The series $\omega_{n}^{(g)}(z_{1},\ldots,z_{n})$ is a well-defined power series in $z_{1}^{-1},\ldots,z_{n}^{-1}$ of the form
$\omega_{n}^{(g)}(z_{1},\ldots,z_{n})=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}\frac{\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{z_{1}^{\ell_{1}+1}\cdots z_{n}^{\ell_{n}+1}}.$ (10)
Its coefficients are related to the generating functions of maps with rooted tight boundaries by
$\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}R^{(\ell_{1}+\cdots+\ell_{n})/2}.$ (11)
Proof.
Observe that
$x(z)^{-1}=\frac{z^{-1}}{R^{1/2}+Sz^{-1}+R^{1/2}z^{-2}}=\frac{z^{-1}}{R^{1/2}}-\frac{Sz^{-2}}{R}+\cdots$ (12)
can be seen as a formal power series in $z^{-1}$ without constant coefficient, so that for any $\ell$, $x^{\prime}(z)/x(z)^{\ell+1}$ is a series containing only powers of $z^{-1}$ larger than or equal to $\ell+1$. This shows that
$\omega_{n}^{(g)}(z_{1},\ldots,z_{n})=\sum_{L_{1},\ldots,L_{n}\geq 1}F^{(g)}_{L_{1},\ldots,L_{n}}\frac{x^{\prime}(z_{1})}{x(z_{1})^{L_{1}+1}}\cdots\frac{x^{\prime}(z_{n})}{x(z_{n})^{L_{n}+1}}$ (13)
is a well-defined power series in $z_{1}^{-1},\ldots,z_{n}^{-1}$ of the form (10): the only terms of (13) contributing to the coefficient $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ of $z_{1}^{-\ell_{1}-1}\cdots z_{n}^{-\ell_{n}-1}$ are those with $1\leq L_{i}\leq\ell_{i}$, for all $i$.
Now we claim that, for any formal Laurent series $F(x)$ in $x^{-1}$, we have the identity
$[x^{-1}]F(x)=[z^{-1}]F(x(z))x^{\prime}(z).$ (14)
Indeed, by linearity it suffices to check this identity for $F(x)=x^{k}$, $k\in\mathbb{Z}$: both sides are equal to $1$ for $k=-1$, and vanish for $k\neq-1$ (since $x(z)^{k}x^{\prime}(z)=\frac{d}{dz}\frac{x(z)^{k+1}}{k+1}$ and a derivative contains no monomial in $z^{-1}$). Applying the identity for each variable of the multivariate formal Laurent series $x_{1}^{L_{1}}\cdots x_{n}^{L_{n}}W_{n}^{(g)}(x_{1},\ldots,x_{n})$, we obtain
$\begin{split}F^{(g)}_{L_{1},\ldots,L_{n}}&=[x_{1}^{-1}\cdots x_{n}^{-1}]x_{1}^{L_{1}}\cdots x_{n}^{L_{n}}W_{n}^{(g)}(x_{1},\ldots,x_{n})\\ &=[z_{1}^{-1}\cdots z_{n}^{-1}]x(z_{1})^{L_{1}}\cdots x(z_{n})^{L_{n}}\omega_{n}^{(g)}(z_{1},\ldots,z_{n})\\ &=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}\left([z_{1}^{\ell_{1}}]x(z_{1})^{L_{1}}\right)\cdots\left([z_{n}^{\ell_{n}}]x(z_{n})^{L_{n}}\right)\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}\\ &=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}R^{(\ell_{1}+\cdots+\ell_{n})/2}\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}\end{split}$ (15)
where $A_{L,\ell}$ is as in (6). Comparing with (5), and noting that the matrix $(A_{L,\ell})_{L,\ell\geq 1}$ is unitriangular hence invertible, we get the wanted equality (11). ∎
Remark 2.6.
As we have $x(z)=x(z^{-1})$ and $x^{\prime}(z)=-x^{\prime}(z^{-1})/z^{2}$, we may alternatively view $\omega_{n}^{(g)}(z_{1},\ldots,z_{n})$ as a formal power series in $z_{1},\ldots,z_{n}$ with expansion
$\omega_{n}^{(g)}(z_{1},\ldots,z_{n})=(-1)^{n}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}R^{-(\ell_{1}+\cdots+\ell_{n})/2}z_{1}^{\ell_{1}-1}\cdots z_{n}^{\ell_{n}-1}.$ (16)
As we will see below, $\omega_{n}^{(g)}(z_{1},\ldots,z_{n})$ is actually a rational function of $z_{1},\ldots,z_{n}$, and (10) and (16) are two different expansions of it (at infinity and at zero, respectively).
2.3 Allowing boundary-vertices
Theorem 2.2 deals with the case where all boundaries of the maps at hand are faces. There is a more general statement that also deals with both boundary-faces and boundary-vertices. Let $M_{0}$ be a map of genus $g$ with $n=m+s$ distinguished tight boundaries, the first $m$ of which are rooted faces $f_{1}^{0},\ldots,f_{m}^{0}$, the remaining $s$ being vertices, for some $m,s\geq 0$. Let $M_{i},1\leq i\leq m$ be a brassband with respective boundary-faces $f_{i},F_{i}$, where $f_{i}$ denotes the mouthpiece. We assume that the lengths of $f_{i}$ and of $f^{0}_{i}$ are equal for $1\leq i\leq m$. We let $M=\Phi(M_{0},M_{1},\ldots,M_{m})$ be the map obtained as above by identifying the boundaries of $f_{i}$ and $f_{i}^{0}$ for $1\leq i\leq m$. Then the map $M$ has $m$ rooted boundary-faces and $s$ boundary-vertices, inherited from $M_{0}$.
Theorem 2.7.
Let $g,m,s$ be nonnegative integers such that $n=m+s\geq 1$ and $\chi=2-2g-n\leq 0$. Let $L_{1},\ell_{1},\ldots,L_{m},\ell_{m}$ be fixed positive integers. Then, the mapping $\Phi$ is a bijection between
  • sequences $(M_{0},M_{1},\ldots,M_{m})$, where $M_{0}$ is a map of genus $g$ with $n$ distinguished tight boundaries, the first $m$ of which are rooted boundary-faces of lengths $\ell_{i},1\leq i\leq m$, the remaining $s$ being boundary-vertices, and for $1\leq i\leq m$, $M_{i}$ is a trumpet with mouthpiece length $\ell_{i}$ and whose other external face has length $L_{i}$, and
  • maps $M$ of genus $g$ with $n$ distinguished boundaries, the first $m$ of which are rooted boundary-faces $F_{1},\ldots,F_{m}$ of lengths $L_{1},\ldots,L_{m}$, the remaining $s$ being boundary-vertices, and such that for $1\leq i\leq m$, the minimal length of a closed path homotopic to $F_{i}$ is equal to $\ell_{i}$.
The justification for this statement is exactly the same as for Theorem 2.2. The inverse bijection consists in cutting the map $M$ along the outermost minimal closed paths that are homotopic to the boundary-faces, considering both boundary-faces and boundary-vertices as punctures. In particular, the boundary-vertices are never included in the trumpets $M_{1},\ldots,M_{m}$ that are cut away.
We now consider the enumerative consequences of Theorem 2.7. In order to put boundary-faces and boundary-vertices on a same footing, we denote by $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ the generating function of maps of genus $g$ with $n$ unrooted tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$. It is related to the generating function $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ considered above by
$\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\ell_{1}\cdots\ell_{n}\,T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ (17)
as there are $\ell_{1}\cdots\ell_{n}$ ways to root the boundaries, up to the symmetry factor corresponding to the inverse of the automorphism group order [^3] Note that the objects enumerated by the generating series $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ are not rooted, and therefore can have nontrivial automorphism groups. This is why we have to weigh them by the inverse of the automorphism group order, as discussed in the introduction.. In particular, there are zero ways if one $\ell_{i}$ vanishes: as said in Section 1.2, a boundary-vertex cannot be rooted. Thus, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ contains a priori more information than $\hat{T}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. In particular, for given integers $m,s\geq 0$ not both equal to 00, and positive integers $\ell_{1},\ldots,\ell_{m}$, the generating function of maps with $m$ rooted tight boundary-faces of lengths $\ell_{1},\ldots,\ell_{m}$ and with $s$ boundary-vertices is given by $(\ell_{1}\cdots\ell_{m})T^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$, with $s$ zero indices. With this remark at hand, we can now state the generating function counterpart to Theorem 2.7 as follows.
Corollary 2.8.
Let $g,m,s$ be nonnegative integers such that $n=m+s\geq 1$ and $\chi=2-2g-n\leq 0$. Let $L_{1},\ldots,L_{m}$ be fixed positive integers. Then, we have
$\frac{\partial^{s}F^{(g)}_{L_{1},\ldots,L_{m}}}{\partial t^{s}}=\sum_{\ell_{1},\ldots,\ell_{m}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{m},\ell_{m}}\,(\ell_{1}\cdots\ell_{m})\,T^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}\,,$ (18)
where there are $s$ indices equal to 00 in the last term, and where we reuse the notations from Corollary 2.3. For $m=0$, the relation reads
$\frac{\partial^{s}F^{(g)}}{\partial t^{s}}=T^{(g)}_{0,\ldots,0}\,,$ (19)
with $s$ indices equal to 00, and where $F^{(g)}$ is the generating function of maps of genus $g$ without boundaries.
For $g,m,s$ as above, we introduce the following generalization of the grand generating function (7):
$W^{(g)}_{m,s}(x_{1},\ldots,x_{m})$ $:=\sum_{L_{1},\ldots,L_{m}}\frac{\partial^{s}F^{(g)}_{L_{1},\ldots,L_{m}}}{\partial t^{s}}\cdot\frac{1}{x_{1}^{L_{1}+1}\cdots x_{m}^{L_{m}+1}}$ (20)
$=\frac{\partial^{s}}{\partial t^{s}}W_{m}^{(g)}(x_{1},\ldots,x_{m})\,,$ (21)
and, assuming further that $\chi<0$, the quantity
$\omega^{(g)}_{m,s}(z_{1},\ldots,z_{m}):=W_{m,s}^{(g)}(x(z_{1}),\ldots,x(z_{m}))x^{\prime}(z_{1})\ldots x^{\prime}(z_{m})\,.$ (22)
For $m=0$, we have $\omega^{(g)}_{0,s}=W_{0,s}^{(g)}={\partial^{s}F^{(g)}}/{\partial t^{s}}$. Corollary 2.8 admits the following consequence, analog to Theorem 2.5, whose proof is adapted from the latter in a straightforward way.
Theorem 2.9.
For integers $\ell_{1},\ldots,\ell_{n}$, define
$\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}:=T^{(g)}_{\ell_{1},\ldots,\ell_{n}}R^{-(\ell_{1}+\cdots+\ell_{n})/2}$ (23)
with $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ the generating function of maps with unrooted tight boundaries defined above. Then, we have the expansion
$\omega_{m,s}^{(g)}(z_{1},\ldots,z_{m})=\sum_{\ell_{1},\ldots,\ell_{m}\geq 1}\frac{\ell_{1}\cdots\ell_{m}\tau^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}}{z_{1}^{\ell_{1}+1}\cdots z_{m}^{\ell_{m}+1}}$ (24)
where there are $s$ trailing zeros in the indices of $\tau^{(g)}$. For $m=0$, we have $\omega^{(g)}_{0,s}=\tau^{(g)}_{0,\ldots,0}=T^{(g)}_{0,\ldots,0}$.
Theorems 2.5 and 2.9 provide combinatorial interpretations for the coefficients of $\omega^{(g)}_{n}$ and $\omega^{(g)}_{m,s}$ in terms of generating series of maps with tight boundaries. However, they do not provide any information about the structural properties of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. This is the goal of the upcoming Section 4, where we will see that these are quasi-polynomials in the variables $\ell_{1},\ldots,\ell_{n}$, and that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{m},0,\ldots,0}$ (with $s$ zero indices) indeed corresponds to the evaluation of this quasi-polynomial function of $n=m+s$ variables when the last $s$ ones are set to 00.
2.4 A formula for the series of planar bipartite maps with tight boundaries
In this section, we give an explicit expression for the generating function $T^{(0)}_{\ell_{1},\ldots,\ell_{n}}$ of planar maps with $n$ unrooted tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$, in the bipartite case where the $\ell_{i}$ are even and where the weights $t_{1},t_{3},t_{5},\ldots$ for inner faces of odd degrees are set to zero. To emphasize this latter restriction we will use the notation $\left.\right|_{\mathrm{bip}}$ in the following equations. Recall that the basic generating function $R$ is here determined by (4).
Our input is the Collet-Fusy formula [CoFu12, Theorem 1.1] for the generating function of planar bipartite maps with $n$ rooted, non necessarily tight, boundaries of prescribed lengths. In our present notations, it reads
$\left.F^{(0)}_{2L_{1},\ldots,2L_{n}}\right|_{\mathrm{bip}}=\frac{L_{1}\cdots L_{n}}{L_{1}+\cdots+L_{n}}\prod_{i=1}^{n}\binom{2L_{i}}{L_{i}}\frac{\partial^{n-2}}{\partial t^{n-2}}R^{L_{1}+\ldots+L_{n}}$ (25)
for any $n\geq 1$ and positive integers $L_{1},\ldots,L_{n}$. Note that the formula makes sense for $n=1$, upon understanding that $\frac{\partial^{-1}}{\partial t^{-1}}$ means integrating over $t$.
We will deduce from the Collet-Fusy formula an expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\rvert_{\mathrm{bip}}$, using Corollary 2.8. Interestingly, our expression involves a certain family of polynomials that we encountered previously in [polytightmaps]. Namely, for any integers $k\geq 0$ and $n\geq 1$, let us consider the multivariate polynomial in the variables $\ell_{1},\ldots,\ell_{n}$
$p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n}):=\sum_{k_{1},k_{2},\ldots,k_{n}\geq 0\atop k_{1}+k_{2}+\cdots+k_{n}=k}p_{k_{1}}(\ell_{1})q_{k_{2}}(\ell_{2})\cdots q_{k_{n}}(\ell_{n})\,,$ (26)
where, in the right-hand side, all factors except the first one are $q_{k_{i}}$’s, and where the univariate polynomials $p_{k}$ and $q_{k}$ are defined by
$\begin{split}&p_{k}(\ell):=\frac{1}{(k!)^{2}}\,\prod_{i=1}^{k}\left(\ell^{2}-i^{2}\right)=\binom{\ell-1}{k}\binom{\ell+k}{k}\\ &q_{k}(\ell):=\frac{1}{(k!)^{2}}\,\prod_{i=0}^{k-1}\left(\ell^{2}-i^{2}\right)=\binom{\ell}{k}\binom{\ell+k-1}{k}\\ \end{split}$ (27)
with the convention $p_{0}(\ell)=q_{0}(\ell)=1$, and with $\binom{x}{k}=x(x-1)\cdots(x-k+1)/k!$ viewed as a polynomial in $x$. Clearly, $p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n})$ is a polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, as discussed in [polytightmaps, Proposition 2.1], it is a symmetric polynomial and it satisfies the consistency relation
$p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n},0)=p_{k}(\ell_{1},\ell_{2},\ldots,\ell_{n})\,.$ (28)
For $k=1,2,3$, we have
$\begin{split}p_{1}(\ell_{1},\ldots,\ell_{n})&=\left(\sum_{i=1}^{n}\ell_{i}^{2}\right)-1,\\ p_{2}(\ell_{1},\ldots,\ell_{n})&=\frac{1}{4}\left(\sum_{i=1}^{n}\ell_{i}^{4}\right)+\sum_{i<j}\ell_{i}^{2}\ell_{j}^{2}-\frac{5}{4}\left(\sum_{i=1}^{n}\ell_{i}^{2}\right)+1,\\ p_{3}(\ell_{1},\ldots,\ell_{n})&=\frac{1}{36}\left(\sum_{i=1}^{n}\ell_{i}^{6}\right)+\frac{1}{4}\left(\sum_{i\neq j}\ell_{i}^{4}\ell_{j}^{2}\right)+\sum_{i<j<h}\ell_{i}^{2}\ell_{j}^{2}\ell_{h}^{2}\\ &\qquad-\frac{7}{18}\left(\sum_{i=1}^{n}\ell_{i}^{4}\right)-\frac{3}{2}\left(\sum_{i<j}\ell_{i}^{2}\ell_{j}^{2}\right)+\frac{49}{36}\left(\sum_{i=1}^{n}\ell_{i}^{2}\right)-1.\end{split}$ (29)
We will also need the so-called partial exponential Bell polynomials, defined for any nonnegative integers $n\geq k\geq 1$ by
$B_{n,k}(r_{1},r_{2},\ldots,r_{n-k+1}):=\\ \sum_{\begin{subarray}{c}j_{1},j_{2},\ldots,j_{n-k+1}\geq 0\\ j_{1}+j_{2}+\cdots+j_{n-k+1}=k\\ j_{1}+2j_{2}+\cdots+(n-k+1)j_{n-k+1}=n\end{subarray}}\!\!\!\!\!\!\!\!\frac{n!}{j_{1}!j_{2}!\cdots j_{n-k+1}!}\left(\frac{r_{1}}{1!}\right)^{j_{1}}\left(\frac{r_{2}}{2!}\right)^{j_{2}}\cdots\left(\frac{r_{n-k+1}}{(n-k+1)!}\right)^{j_{n-k+1}}.$ (30)
Combinatorially, $B_{n,k}(r_{1},r_{2},\ldots)$ is the generating series of partitions of an $n$-element set into $k$ blocks, where we attach a weight $r_{i}$ to eack block of size $i$. Table 1 lists the first few Bell polynomials. They appear in the following classical formula:
Proposition 2.10 (Faà di Bruno’s formula).
Let $f,g$ be sufficiently smooth functions of one variable, and let $n$ be a positive integer. Then, we have
$\frac{d^{n}}{dt^{n}}(f\circ g)(t)=\sum_{k=1}^{n}f^{(k)}(g(t))B_{n,k}\left(g^{\prime}(t),g^{\prime\prime}(t),\ldots,g^{(n-k+1)}(t)\right)\,.$ (31)
$n$ $k$ $1$ $2$ $3$ $4$
$1$ $r_{1}$
$2$ $r_{2}$ $r_{1}^{2}$
$3$ $r_{3}$ $3r_{1}r_{2}$ $r_{1}^{3}$
$4$ $r_{4}$ $4r_{1}r_{3}+3r_{2}^{2}$ $6r_{1}^{2}r_{2}$ $r_{1}^{4}$
Table 1: The first few partial exponential Bell polynomials $B_{n,k}(r_{1},r_{2},\ldots)$.
See for instance the discussion in [FlSe09, Example III.24]. We may now state the main result of this section:
Theorem 2.11.
For $n\geq 3$ and $\ell_{1},\ldots,\ell_{n}$ nonnegative integers, the generating function of bipartite planar maps with $n$ unrooted tight boundaries of lengths $2\ell_{1},\ldots,2\ell_{n}$ is equal to
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}}\sum_{k=0}^{n-3}k!p_{k}(\ell_{1},\ldots,\ell_{n})b_{n-2,k+1}+\frac{(-1)^{n}(n-3)!}{t^{n-2}}\delta_{\ell_{1}+\cdots+\ell_{n},0}$ (32)
with $R$ given by (4), and where we introduce the shorthand notation
$b_{n,k}:=R^{-k}B_{n,k}(R^{\prime},R^{\prime\prime},\ldots)=B_{n,k}\left(\frac{R^{\prime}}{R},\frac{R^{\prime\prime}}{R},\ldots\right)$ (33)
with $R^{\prime},R^{\prime\prime},\ldots$ the successive derivatives of $R$ with respect to the vertex weight $t$.
For $n=3$ we recover $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3}}\rvert_{\mathrm{bip}}=R^{\ell_{1}+\ell_{2}+\ell_{3}-1}R^{\prime}-t^{-1}\delta_{\ell_{1}+\ell_{2}+\ell_{3},0}$ consistently with [triskell, Theorem 1.1], while for $n=4$ we find
$\left.T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\ell_{2}+\ell_{3}+\ell_{4}}\left(\left(\ell_{1}^{2}+\ell_{2}^{2}+\ell_{3}^{2}+\ell_{4}^{2}-1\right)\frac{R^{\prime 2}}{R^{2}}+\frac{R^{\prime\prime}}{R}\right)+\frac{\delta_{\ell_{1}+\cdots+\ell_{4},0}}{t^{2}}.$ (34)
Longer expressions for $n=5,6$ may be written straightforwardly from (29) and Table 1.
Remark 2.12.
When setting the weights $t_{2},t_{4},\ldots$ for inner faces to zero, we have $R=t$, $R^{\prime}=1$ and $R^{(j)}=0$ for $j\geq 2$. Thus, in this case we have $b_{n,k}=t^{-k}\delta_{n,k}$, hence $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ reduces to $(n-3)!p_{n-3}(\ell_{1},\ldots,\ell_{n})t^{\ell_{1}+\ldots+\ell_{n}-n+2}$ if at least one $\ell_{i}$ is non-zero, and vanishes otherwise since $p_{n-3}(0,\ldots,0)=(-1)^{n-3}$ so the rightmost term in (32) cancels the result. This is consistent with [polytightmaps, Theorem 2.3] since the tight maps considered in this reference are nothing but maps with tight boundaries and without inner faces. Note that the exponent of $t$ is consistent with Euler’s relation.
Proof of Theorem 2.11.
Let us first consider the case where all $\ell_{i}$ are zero: then $T^{(0)}_{0,\ldots,0}$ is simply the generating function of bipartite planar maps with $n\geq 3$ marked vertices. For $n=3$ it is equal to $\frac{\partial}{\partial t}\ln(R/t)$, see e.g. [triskell, Appendix A], and for larger $n$ we simply have to take more derivatives with respect to $t$, namely
$\left.T^{(0)}_{0,\ldots,0}\right|_{\mathrm{bip}}=\frac{\partial^{n-2}}{\partial t^{n-2}}\ln(R/t)=\frac{\partial^{n-2}}{\partial t^{n-2}}\ln R+\frac{(-1)^{n}(n-3)!}{t^{n-2}}\,.$ (35)
Now, by applying Faà di Bruno’s formula, we get
$\left.T^{(0)}_{0,\ldots,0}\right|_{\mathrm{bip}}=\sum_{k=1}^{n-2}(-1)^{k-1}(k-1)!R^{-k}B_{n-2,k}(R^{\prime},R^{\prime\prime},\ldots)+\frac{(-1)^{n}(n-3)!}{t^{n-2}}\\ =\sum_{k=0}^{n-3}(-1)^{k}k!b_{n-2,k+1}+\frac{(-1)^{n}(n-3)!}{t^{n-2}}.$ (36)
This gives the wanted formula for $\ell_{1}=\cdots=\ell_{n}=0$ since, by (27) and (28), we have $p_{k}(0,\ldots,0)=p_{k}(0)=(-1)^{k}$.
Let us now assume that at least one $\ell_{i}$ is non-zero, and let $m\geq 1$ be the number of non-zero $\ell_{i}$, and $s=n-m$ the number of zero $\ell_{i}$. Without loss of generality, we can assume that $\ell_{1},\ldots,\ell_{m}>0$ and $\ell_{i}=0$ for $i>m$. We note that, in the bipartite case, the trumpet generating function of Proposition 2.4 vanishes whenever its indices are of different parities (since $S=0$), and reads for even indices
$A_{2L,2\ell}=\binom{2L}{L-\ell}R^{L-\ell}\,.$ (37)
Thus, Corollary 2.8 reduces to
$\frac{\partial^{s}}{\partial t^{s}}\left.F^{(0)}_{2L_{1},\ldots,2L_{m}}\right|_{\mathrm{bip}}=\sum_{\ell_{1},\ldots,\ell_{m}\geq 1}(2\ell_{1})A_{2L_{1},2\ell_{1}}\cdots(2\ell_{m})A_{2L_{m},2\ell_{m}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{m},0,\ldots,0}\right|_{\mathrm{bip}}$ (38)
(where we append $s$ zeros after $2\ell_{m}$). On the other hand, by the Collet-Fusy formula and by Faà di Bruno’s formula, we have
$\frac{\partial^{s}}{\partial t^{s}}\left.F^{(0)}_{2L_{1},\ldots,2L_{m}}\right|_{\mathrm{bip}}=\frac{L_{1}\cdots L_{m}}{L_{1}+\cdots+L_{m}}\binom{2L_{1}}{L_{1}}\cdots\binom{2L_{m}}{L_{m}}\frac{\partial^{n-2}}{\partial t^{n-2}}R^{L_{1}+\cdots+L_{m}}\\ =\sum_{k=1}^{n-2}L_{1}\cdots L_{m}(L_{1}+\cdots+L_{m}-1)_{k-1}\binom{2L_{1}}{L_{1}}\cdots\binom{2L_{m}}{L_{m}}R^{L_{1}+\cdots+L_{m}}b_{n-2,k}$ (39)
where $(L)_{k}:=L(L-1)\cdots(L-k+1)$ denotes the falling factorial. We will rewrite this expression in the same form as the right-hand side of (38), so as to identify $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{m},0,\ldots,0}\right|_{\mathrm{bip}}$. To this end, we proceed as in [polytightmaps, Section 3.1] and use the Chu-Vandermonde identity to write
$(L_{1}+L_{2}+\cdots+L_{m}-1)_{k-1}=(k-1)!\sum_{\begin{subarray}{c}k_{1},k_{2},\ldots,k_{m}\geq 0\\ k_{1}+k_{2}+\cdots+k_{m}=k-1\end{subarray}}\binom{L_{1}-1}{k_{1}}\binom{L_{2}}{k_{2}}\cdots\binom{L_{m}}{k_{m}}.$ (40)
Plugging into (39) we obtain
$\frac{\partial^{s}}{\partial t^{s}}\left.F^{(0)}_{2L_{1},\ldots,2L_{m}}\right|_{\mathrm{bip}}=\sum_{k=1}^{n-2}(k-1)!R^{L_{1}+\cdots+L_{m}}b_{n-2,k}\times\\ \sum_{\underset{k_{1}+\cdots+k_{m}=k-1}{k_{1},\ldots,k_{m}\geq 0}}L_{1}\binom{L_{1}-1}{k_{1}}\binom{2L_{1}}{L_{1}}L_{2}\binom{L_{2}}{k_{2}}\binom{2L_{2}}{L_{2}}\cdots L_{m}\binom{L_{m}}{k_{m}}\binom{2L_{m}}{L_{m}}\,.$ (41)
Now, by [polytightmaps, Lemma 3.1] and by (37) [^4] Beware that the notation $A_{\ell,m}$ used in [polytightmaps] differs from that of the present paper., we have for $L>0$
$\begin{split}L\binom{L-1}{k}\binom{2L}{L}&=\sum_{\ell=1}^{L}(2\ell)\binom{2L}{L-\ell}p_{k}(\ell)=\sum_{\ell=1}^{L}(2\ell)A_{2L,2\ell}R^{\ell-L}p_{k}(\ell)\,,\\ L\binom{L}{k}\binom{2L}{L}&=\sum_{\ell=1}^{L}(2\ell)\binom{2L}{L-\ell}q_{k}(\ell)=\sum_{\ell=1}^{L}(2\ell)A_{2L,2\ell}R^{\ell-L}q_{k}(\ell)\,.\end{split}$ (42)
Therefore, by the definition (26) of $p_{k}$, we may rewrite (41) as
$\frac{\partial^{s}}{\partial t^{s}}\left.F^{(0)}_{2L_{1},\ldots,2L_{m}}\right|_{\mathrm{bip}}=\sum_{\ell_{1},\ldots,\ell_{m}\geq 1}(2\ell_{1})A_{2L_{1},2\ell_{1}}\cdots(2\ell_{m})A_{2L_{m},2\ell_{m}}\\ \times R^{\ell_{1}+\ldots+\ell_{m}}\sum_{k=1}^{n-2}(k-1)!p_{k-1}(\ell_{1},\ldots,\ell_{m})b_{n-2,k}\,,$ (43)
and by (28), we may replace $p_{k-1}(\ell_{1},\ldots,\ell_{m})$ by $p_{k-1}(\ell_{1},\ldots,\ell_{m},0,\ldots,0)$, in order to have $n$ variables. Comparing this expression with (38), using the invertibility of the unitriangular matrix $(A_{2L,2\ell})_{L,\ell\geq 1}$ and doing a change of variable $k-1\to k$ finally gives the wanted formula (32), since the term $\frac{(-1)^{n}(n-3)!}{t^{n-2}}$ is absent for $s\geq 1$. ∎
A variant of the Collet-Fusy formula (25) holds in the quasi-bipartite case where exactly two of the boundaries have odd lengths, i.e. when exactly two of the $L_{i}$, say $L_{1}$ and $L_{2}$, are half-integers. In this case, the factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ appearing in the right-hand side should be replaced by $4\binom{2L_{1}-1}{L_{1}-1/2}\binom{2L_{2}-1}{L_{2}-1/2}$. Correspondingly, we have the following analogue of Theorem 2.11:
Proposition 2.13.
For $n\geq 3$, $\ell_{1},\ell_{2}$ positive half-integers and $\ell_{3},\ldots,\ell_{n}$ nonnegative integers, the generating function of quasi-bipartite planar maps with $n$ unrooted tight boundaries of lengths $2\ell_{1},\ldots,2\ell_{n}$ is equal to
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}}\sum_{k=0}^{n-3}k!\tilde{p}_{k}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})b_{n-2,k+1}$ (44)
where, following [polytightmaps, Section 2.1.3], we define
$\tilde{p}_{k}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n}):=\sum_{k_{1},k_{2},\ldots,k_{n}\geq 0\atop k_{1}+k_{2}+\cdots+k_{n}=k}\tilde{p}_{k_{1}}(\ell_{1})\tilde{p}_{k_{2}}(\ell_{2})q_{k_{3}}(\ell_{3})\cdots q_{k_{n}}(\ell_{n})$ (45)
with $q_{k}$ as in (27) and with
$\tilde{p}_{k}(\ell):=\frac{1}{(k!)^{2}}\,\prod_{i=1}^{k}\left(\ell^{2}-\left(i-\frac{1}{2}\right)^{2}\right)=\binom{\ell-\frac{1}{2}}{k}\binom{\ell+k-\frac{1}{2}}{k}.$ (46)
Proof.
Starting with the quasi-bipartite variant of the Collet-Fusy formula, we proceed as in the proof of Theorem 2.11. Using Faà di Bruno’s formula, we obtain the expression (39) with the binomial factors $\binom{2L_{1}}{L_{1}}\binom{2L_{2}}{L_{2}}$ modified as discussed above. Now, we expand the falling factorial as
$(L_{1}+\cdots+L_{m}-1)_{k-1}=(k-1)!\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 0\\ k_{1}+\cdots+k_{m}=k-1\end{subarray}}\binom{L_{1}-\frac{1}{2}}{k_{1}}\binom{L_{2}-\frac{1}{2}}{k_{2}}\binom{L_{3}}{k_{3}}\cdots\binom{L_{m}}{k_{m}}.$ (47)
which leads to a variant of (41) in which the factors $\binom{L_{1}-1}{k_{1}}\binom{2L_{1}}{L_{1}}\binom{L_{2}}{k_{2}}\binom{2L_{2}}{L_{2}}$ are replaced by $4\binom{L_{1}-1/2}{k_{1}}\binom{2L_{1}-1}{L_{1}-1/2}\binom{L_{2}-1/2}{k_{2}}\binom{2L_{2}-1}{L_{2}-1/2}$. By [polytightmaps, Equation (3.9)], we have for $L$ a positive half-integer
$2L\binom{L-\frac{1}{2}}{k}\binom{2L-1}{L-\frac{1}{2}}=\sum_{\ell=\frac{1}{2}}^{L}(2\ell)\binom{2L}{L-\ell}\tilde{p}_{k}(\ell)=\sum_{\ell=\frac{1}{2}}^{L}(2\ell)A_{2L,2\ell}R^{\ell-L}\tilde{p}_{k}(\ell)$ (48)
where the sum runs over the positive half-integers $\ell$ between $\frac{1}{2}$ and $L$. Using again the second line of (42) to expand the remaining binomial factors, we arrive at a variant of (43) with $p_{k-1}$ replaced by $\tilde{p}_{k-1}$, giving the wanted result. ∎
Let us consider the case of four boundaries. We have $\tilde{p}_{1}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=\ell_{1}^{2}+\cdots+\ell_{4}^{2}-\frac{1}{2}$ and $\tilde{p}_{0}(\ell_{1},\ell_{2};\ell_{3},\ell_{4})=1$, hence from Proposition 2.13 we see that, in the quasi-bipartite case, the term $-1$ appearing in the expression (34) should be replaced by $-\frac{1}{2}$. In contrast, we found after a lengthy computation (which we omit here) that (34) holds as is when all $\ell_{i}$ are half-integers, i.e. when all four boundaries have odd lengths. This is consistent with the discussion of the lattice count polynomial $N_{0,4}$ at the end of [Norbury2010, Section 1], since $T^{(0)}_{2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4}}$ reduces to $N_{0,4}(2\ell_{1},2\ell_{2},2\ell_{3},2\ell_{4})$ when we set the weights for inner faces to zero.
3 Recursion relations for series of maps with tight boundaries
In this section we give a number of recursion relations for the generating function $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, as obtained by adding an extra tight boundary to a map with pre-existing boundaries. We will need the following proposition, proved bijectively in Appendix A:
Proposition 3.1.
Let $\ell_{1}$, $\ell_{2}$ and $\ell_{3}$ be nonnegative integers and let $T^{(0)}_{\ell_{1},\ell_{2}|\ell_{3}}$ denote the generating function of planar maps with three labeled distinct tight unrooted boundaries of lengths $\ell_{1},\ell_{2},\ell_{3}$, the third being strictly tight, where we attach a weight $t$ per vertex different from a boundary-vertex and not incident to the third boundary (and for all $k\geq 1$, a weight $t_{k}$ per inner face of degree $k$). Then, for $\ell_{1}+\ell_{2}>\ell_{3}$, we have
$T^{(0)}_{\ell_{1},\ell_{2}|\ell_{3}}=\begin{cases}R^{(\ell_{1}+\ell_{2}-\ell_{3})/2-1}\displaystyle{\frac{\partial R}{\partial t}}&\text{if $\ell_{1}+\ell_{2}-\ell_{3}$ is even,}\\ \\ R^{(\ell_{1}+\ell_{2}-\ell_{3}-1)/2}\displaystyle{\frac{\partial S}{\partial t}}&\text{if $\ell_{1}+\ell_{2}-\ell_{3}$ is odd.}\end{cases}$ (49)
3.1 Adding an extra boundary-vertex
Marking a vertex in a trumpet.
Take $L,\ell$ positive integers. Recall that $A_{L,\ell}$ is the generating function for trumpets with rooted boundary of length $L$ and mouthpiece of length $\ell$. The quantity $\frac{\partial A_{L,\ell}}{\partial t}$ is therefore the generating function for such trumpets with a marked vertex which is not incident to the mouthpiece (recall that vertices of the mouthpiece do not receive the weight $t$). It has the following expression:
Proposition 3.2.
We have
$\begin{split}\frac{\partial A_{L,\ell}}{\partial t}&=\sum_{m>\ell}A_{L,m}\,m\,T^{(0)}_{m,0|\ell}\\ &=\sum_{m=\ell+1\atop m=\ell+1\!\!\!\!\mod 2}^{L}A_{L,m}\,m\,R^{\frac{m-\ell-1}{2}}\frac{\partial S}{\partial t}+\sum_{m=\ell+2\atop m=\ell+2\!\!\!\!\mod 2}^{L}A_{L,m}\,m\,R^{\frac{m-\ell}{2}-1}\frac{\partial R}{\partial t}\,.\end{split}$ (50)
Combinatorial proof.
Considering the marked vertex as a boundary-vertex, we obtain a planar map with three boundaries which me may decompose as follows: consider the outermost minimal closed path homotopic to the boundary of length $L$. Cutting along this path produces a trumpet whose mouthpiece has length $m>\ell$, see Figure 3. The remaining map with three tight boundaries is precisely a map enumerated by $T^{(0)}_{m,0|\ell}$, endowed canonically with one of its $m$ rootings (inherited from the rooting of the boundary face of length $L$). ∎
Figure 3: An illustration of the decomposition of a trumpet with mouthpiece of length $\ell$ endowed with an extra marked vertex, viewed as a boundary vertex. It results in two pieces: a trumpet with mouthpiece of length $m>\ell$ and a map with three boundaries counted by $m\,T^{(0)}_{m,0|\ell}$.
It is interesting to note that the expression in the second line in Eq. (50) may be obtained in a purely computational way as follows:
Computational proof.
From the explicit expression (6) of $A_{L,\ell}$, we have
$\begin{split}\frac{\partial A_{L,\ell}}{\partial t}&=\frac{\partial\ }{\partial t}\left([z^{\ell}]\left(z+S+\frac{R}{z}\right)^{L}\right)\\ &=[z^{\ell}]\ L\left(z+S+\frac{R}{z}\right)^{L-1}\left(\frac{\partial S}{\partial t}+\frac{1}{z}\frac{\partial R}{\partial t}\right)\\ &=L\,A_{L-1,\ell}\frac{\partial S}{\partial t}+L\,A_{L-1,\ell+1}\frac{\partial R}{\partial t}.\end{split}$ (51)
To obtain the second line of (50), it is enough to prove the equality
$L\,A_{L-1,\ell}=\sum_{p\geq 0}(\ell+2p+1)R^{p}A_{L,\ell+2p+1}\ ,$ (52)
which itself follows from iterating the identity
$L\,A_{L-1,\ell}=(\ell+1)RA_{L,\ell+1}+R\,L\,A_{L-1,\ell+2}$ (53)
$p$ times until $p$ reaches a value such that $\ell+2p>L-1$. As for this last identity (53), upon moving its last term to the left hand side, it is simply obtained by extracting the $z^{\ell}$ coefficient of the straightforward identity
$\left(1-\frac{R}{z^{2}}\right)\frac{\partial\ }{\partial S}\left(z+S+\frac{R}{z}\right)^{L}=\frac{\partial\ }{\partial z}\left(z+S+\frac{R}{z}\right)^{L}\ .$ (54)
Marking a vertex in a map with tight boundaries.
We start with equation (18) with $m=n$ and $s=0$, which we differentiate with respect to $t$. We get
$\begin{split}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}&=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}\\ &+\sum_{i=1}^{n}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots\frac{\partial A_{L_{i},\ell_{i}}}{\partial t}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .\\ \end{split}$ (55)
From (50), the general term in the final sum over $i$ may be written as
$\begin{split}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&\left(\sum_{m_{i}>\ell_{i}}A_{L_{i},m_{i}}\,m_{i}\,T^{(0)}_{m_{i},0|\ell_{i}}\right)\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{i},\cdots\ell_{n}}\\ =\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots&A_{L_{i},\ell_{i}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{i}\cdots\ell_{n})\left(\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\right)\end{split}$ (56)
where, to go from the first to the second line, we exchanged the dummy summation variables $\ell_{i}$ and $m_{i}$, as well as the order of summation. Comparing with the expression for $\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}$ coming from (18) with $m=n$ and $s=1$ , namely
$\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t}=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}\ ,$ (57)
and using the invertibility of the unitriangular matrix $(A_{L,\ell})_{L,\ell\geq 1}$, we obtain the identity
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}.$ (58)
This identity has the following nice combinatorial interpretation: consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}$, and let $v$ be the boundary-vertex corresponding to the last 00. We claim that there exists at most one $i=1,\ldots,n$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating $v$ and the $i$-th boundary from the others. Indeed, if two such $i_{1}$ and $i_{2}$ existed, we could construct a closed path contradicting the tightness of the $i_{1}$-th or the $i_{2}$-th boundary (see Figure 4). If no such $i$ exists, $v$ can be changed into a regular vertex without affecting the tightness of the other boundaries, and we get a map counted by $\frac{\partial}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Otherwise, if exactly one such $i$ exists, we can consider the “extremal” shortest closed path separating $v$ and the $i$-th boundary from the others. By cutting along this path, we get on the one hand a map counted by $T^{(0)}_{\ell_{i},0|m_{i}}$ (it is a pair of pants with one tight boundary of length $\ell_{i}$, one boundary-vertex and one strictly tight boundary of length $m_{i}$) and on the other hand a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue back these maps together.
Note that the above reasoning also works when $\ell_{i}$ is allowed to take the value 00, in which case the sum over $m_{i}$ vanishes. Putting things together, we arrive at the following:
Proposition 3.3.
Let $g,n$ be nonnegative integers. We have, for $\chi=2-2g-n\leq 0$ and $\ell_{1},\ldots,\ell_{n}$ nonnegative integers,
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},0}&=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},0|m_{i}}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &=\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,.\\ \end{split}$ (59)
Note that, for $n=0$, the relation reads $T^{(g)}_{0}=\partial T^{(g)}_{\varnothing}/\partial t=\partial F^{(g)}/\partial t$, where $T^{(g)}_{\varnothing}=F^{(g)}$ is the generating function of maps of genus $g$ without boundaries.
Figure 4: A sketch of why $m_{i_{1}}<\ell_{i_{1}}$ and $m_{i_{2}}<\ell_{i_{2}}$ are mutually exclusive: indeed, it would imply $\mathcal{L}_{i_{1}}<\ell_{i_{1}}$ or $\mathcal{L}_{i_{2}}<\ell_{i_{2}}$, in contradiction with the fact that the boundary-faces of length $\ell_{1}$ and $\ell_{2}$ are both tight.
3.2 Adding an extra boundary-face
Marking an extra boundary-face in a trumpet.
Take $L,\ell,M$ positive integers. The quantity $M\frac{\partial A_{L,\ell}}{\partial t_{M}}$ is the generating function for trumpets with rooted boundary of length $L$, mouthpiece of length $\ell$ and with an extra marked rooted boundary-face of length $M$.
Proposition 3.4.
We have
$M\frac{\partial A_{L,\ell}}{\partial t_{M}}=\sum_{m>\ell}\sum_{\ell_{0}\geq 1}A_{L,m}\,A_{M,\ell_{0}}\,m\,\ell_{0}\,T^{(0)}_{m,\ell_{0}|\ell}.$ (60)
The proof is in all points similar to that of Proposition 50 and amounts to performing a trumpet decomposition of the map with three boundaries of length $L$, $M$ and $\ell$.
Marking extra boundary-faces via boundary insertion operators.
We now wish to find a recurrence relation relating $T^{(g)}_{m_{1},\ldots,m_{n},m_{n+1}}$ to $T^{(g)}_{m_{1},\ldots,m_{n}}$ for $m_{n+1}$ a positive integer.
For $m$ a positive integer, let us define the tight boundary insertion operator $D_{m}$ by
$D_{m}:=\sum_{M=1}^{m}(A^{-1})_{m,M}\frac{M}{m}\frac{\partial\ }{\partial t_{M}}$ (61)
where $A^{-1}$ is the inverse of the unitriangular matrix $A=(A_{M,m})_{M,m\geq 1}$. It satisfies
$M\frac{\partial\ }{\partial t_{M}}=\sum_{m=1}^{M}mA_{M,m}D_{m}\ .$ (62)
We start with the identity
$L_{n+1}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t_{L_{n+1}}}=F^{(g)}_{L_{1},\ldots,L_{n},L_{n+1}}=\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}\,(\ell_{1}\cdots\ell_{n+1})\,T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}\ .$ (63)
Alternatively, we may write
$\begin{split}L_{n+1}\frac{\partial F^{(g)}_{L_{1},\ldots,L_{n}}}{\partial t_{L_{n+1}}}&=\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})L_{n+1}\frac{\partial T^{(g)}_{\ell_{1},\ldots,\ell_{n}}}{\partial t_{L_{n+1}}}\\ &+\sum_{i=1}^{n}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots\left(L_{n+1}\frac{\partial A_{L_{i},\ell_{i}}}{\partial t_{L_{n+1}}}\right)\cdots A_{L_{n},\ell_{n}}(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .\\ \end{split}$ (64)
From (60), the general term in the final sum over $i$ may be written as
$\begin{split}\sum_{\ell_{1},\ldots,\ell_{n}\geq 1}A_{L_{1},\ell_{1}}\cdots\Bigg{(}\sum_{m_{i}>\ell_{i}}\sum_{\ell_{n+1}\geq 1}A_{L_{i},m_{i}}&\,A_{L_{n+1},\ell_{n+1}}m_{i}\,\ell_{n+1}T^{(0)}_{m_{i},\ell_{n+1}|\ell_{i}}\Bigg{)}\cdots A_{L_{n},\ell_{n}}\\ &\qquad\qquad\qquad\qquad\times(\ell_{1}\cdots\ell_{n})T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\\ =\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}(\ell_{1}\cdots\ell_{n+1})&\left(\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\right)\end{split}$ (65)
where, to go from the first to the second line, we exchanged the dummy summation variables $\ell_{i}$ and $m_{i}$, as well as the order of summation. As for the first term in the right hand side of (64), using (62), it may be rewritten as
$\sum_{\ell_{1},\ldots,\ell_{n+1}\geq 1}A_{L_{1},\ell_{1}}\cdots A_{L_{n+1},\ell_{n+1}}(\ell_{1}\cdots\ell_{n+1})D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}\ .$ (66)
Gathering all the terms in (64), comparing with (63), and noting that we may incorporate Proposition 3.3 upon defining
$D_{0}:=\frac{\partial}{\partial t}\,,$ (67)
we arrive at the following identification:
Proposition 3.5.
Let $g,n$ be nonnegative integers. We have, for $\chi=2-2g-n\leq 0$ and $\ell_{1},\ldots,\ell_{n+1}$ nonnegative integers,
$T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\sum_{m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots\ell_{n}}\ .$ (68)
If $\ell_{n+1}$ is even, this reads
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}&=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,,\\ \end{split}$ (69)
while if $\ell_{n+1}$ is odd, it reads instead
$\begin{split}T^{(g)}_{\ell_{1},\ldots,\ell_{n},\ell_{n+1}}&=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\bigg{(}\sum_{m_{i}=1\atop m_{i}=\ell_{i}-1\!\!\!\!\mod 2}^{\ell_{i}-1}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}-1}\frac{\partial R}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\\ &\hskip 93.89418pt+\sum_{m_{i}=1\atop m_{i}=\ell_{i}-2\!\!\!\!\mod 2}^{\ell_{i}-2}\!\!\!\!\!\!\!\!m_{i}\,R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}-1}{2}}\frac{\partial S}{\partial t}T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}\bigg{)}\,.\\ \end{split}$ (70)
For $n=0$, these formulas read $T^{(g)}_{\ell_{1}}=D_{\ell_{1}}T^{(g)}_{\varnothing}=D_{\ell_{1}}F^{(g)}$.
Again, the identity (68) has a clear combinatorial interpretation. Consider a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$. Then, there exists at most one $i\in\{1,\ldots,n\}$ such that there exists a closed path of length $m_{i}<\ell_{i}$ separating the $i$-th and the $(n+1)$-th boundaries from the others, for a reason similar to that illustrated in Figure 4. If exactly one such $i$ exists, we can consider the extremal shortest separating closed path and cut along it. We get, on the one hand, a map counted by $T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}$, i.e. a pair of pants with two tight boundaries of lengths $\ell_{i}$ and $\ell_{n+1}$, and one strictly tight boundary of length $m_{i}$, and, on the other hand, a map counted by $T^{(g)}_{\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n}}$. There are $m_{i}$ ways to glue back these maps together. If no such $i$ exists, the $(n+1)$-th boundary can be changed into an inner face without affecting the tightness of the other boundaries. We deduce that such maps are counted by $D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, which explains, in retrospect, the name of the operator $D_{m}$.
Remark 3.6.
For $g=0$ and $n=2$, notice that Proposition 3.5 relates the generating function of tight cylinders, given by $T^{(0)}_{\ell_{1},\ell_{2}}=\delta_{\ell_{1},\ell_{2}}R^{\ell_{1}}/\ell_{1}$ for $\ell_{1}>0$, and $T^{(0)}_{0,0}=\ln(R/t)$, with that of tight pairs of pants given by Theorem A.1 below.
3.3 Consistency with the expression for the planar bipartite case
In this section we check that the explicit expression (32) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar bipartite case is consistent with the above relations (59) and (68) describing the addition of an extra boundary-vertex and an extra boundary-face respectively.
Adding an extra boundary-vertex.
The relation (59) of Proposition 3.3 reduces in the planar bipartite case to
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}=\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ (71)
with $R$ given by (4). Let us see how this relation is indeed satisfied by the expression (32) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$.
By (32), we have indeed
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}}\sum_{k=0}^{n-2}k!p_{k}(\ell_{1},\ldots,\ell_{n})b_{n-1,k+1}+\frac{(-1)^{n+1}(n-2)!}{t^{n-1}}\delta_{\ell_{1}+\cdots+\ell_{n},0}$ (72)
while
$\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}\sum_{k=0}^{n-3}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)R^{\prime}b_{n-2,k+1}+\sum_{j\geq 1}R^{(j+1)}b_{n-2,k+1}^{[j]}\right)\\ +\frac{(-1)^{n+1}(n-2)!}{t^{n-1}}\delta_{\ell_{1}+\cdots+\ell_{n},0}\ ,$ (73)
where $R^{(i)}$ denotes the $i$-th derivative of $R$ with respect to $t$ while $b_{n,k}^{[j]}$ is the derivative of the Bell polynomial $B_{n,k}$ with respect to its $j$-th variable, evaluated at $(R^{\prime}/R,R^{\prime\prime}/R,\ldots)$. By the recursion relation
$b_{n-1,k+1}=\frac{R^{\prime}}{R}b_{n-2,k}+\sum_{j\geq 1}\frac{R^{(j+1)}}{R}b_{n-2,k+1}^{[j]}\ ,$ (74)
which is easily derived from the combinatorial interpretation of the Bell polynomials in terms of partitions, we deduce that
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}-\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}R^{\prime}\times\\ \sum_{k=0}^{n-2}k!p_{k}(\ell_{1},\ldots,\ell_{n})\left(b_{n-2,k}-(\ell_{1}+\cdots+\ell_{n}-k-1)b_{n-2,k+1}\right).$ (75)
On the other hand we have
$\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}R^{\prime}\times\\ \sum_{k=0}^{n-2}\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})k!p_{k}(\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n})b_{n-2,k+1}.$ (76)
Comparing (75) and (76), we see that the relation (71) is then a direct consequence of the following recursion relation for $p_{k}(\ell_{1},\ldots,\ell_{n})$:
Proposition 3.7.
(String equation) For nonnegative integers $\ell_{1},\ldots,\ell_{n}$, we have
$(k+1)p_{k+1}(\ell_{1},\ldots,\ell_{n})=(\ell_{1}+\cdots+\ell_{n}-k-1)p_{k}(\ell_{1},\ldots,\ell_{n})\\ +\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})p_{k}(\ell_{1},\ldots,\ell_{i-1},m_{i},\ell_{i+1},\ldots,\ell_{n})\ .$ (77)
This latter recursion is proved in [polytightmaps, Proposition 2.2].
Adding an extra boundary-face of even length.
We now wish to check the consistency of the expression (32) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ with the recursion relation (68) of Proposition 3.5 which, in the planar bipartite case, reduces to
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}.$ (78)
As a prerequisite to prove (78) from (32), we note that $D_{2m}$ satisfies the Leibniz product rule, namely $D_{2m}(XY)=(D_{2m}X)Y+X(D_{2m}Y)$, and moreover that the action of $D_{2m}$ on derivatives of $R$ is given by
$D_{2m}R^{(j)}=T^{(0)}_{2m,2,\!\!\!\!\underbrace{\scriptstyle 0,\ldots,0}_{j+1\text{ times}}}\Big{|}_{\mathrm{bip}}=R^{m+1}\sum_{k=0}^{j}k!q_{k}(m)b_{j+1,k+1}.$ (79)
Here, we may justify the first equality by noting that, in the bipartite case, $R^{(j)}$ can be seen as the generating function of maps with one unrooted boundary-face of length $2$ and $j+1$ boundary-vertices. Since $D_{2m}$ creates an extra tight boundary-face of length $2m$, and since the other boundaries (which have lengths $2$ and 00 only) are automatically tight, we indeed obtain the maps counted by $T^{(0)}_{2m,2,\!\!\!\!\underbrace{\scriptstyle 0,\ldots,0}_{j+1\text{ times}}}\Big{|}_{\mathrm{bip}}$. As for the second equality, it is a particular instance of (32), simplified via the identity $p_{k}(m,1,0,\ldots,0)=p_{k}(m,1)=p_{k}(m)+p_{k-1}(m)=q_{k}(m)$.
By (32), we now have
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}}\sum_{k=0}^{n-2}k!p_{k}(\ell_{1},\ldots,\ell_{n},\ell_{n+1})b_{n-1,k+1}\\ =R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}}\sum_{k^{\prime}\geq 0}\sum_{k^{\prime\prime}\geq 0}(k^{\prime}+k^{\prime\prime})!p_{k^{\prime\prime}}(\ell_{1},\ldots,\ell_{n})q_{k^{\prime}}(\ell_{n+1})b_{n-1,k^{\prime}+k^{\prime\prime}+1}$ (80)
(throughtout this computation, we will leave the summation ranges as implicit as possible, since they are naturally enforced by the vanishing of the summands). On the other hand, we have, using (79),
$D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}-1}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)(D_{2\ell_{n+1}}R)b_{n-2,k+1}+\sum_{j\geq 1}(D_{2\ell_{n+1}}R^{(j)})b_{n-2,k+1}^{[j]}\right)\\ =R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}}\sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\times\hfill\\ \left((\ell_{1}+\cdots+\ell_{n}-k-1)\frac{R^{\prime}}{R}b_{n-2,k+1}+\sum_{j\geq 1}\sum_{k^{\prime}\geq 0}k^{\prime}!q_{k^{\prime}}(\ell_{n+1})b_{j+1,k^{\prime}+1}b_{n-2,k+1}^{[j]}\right)\ .$ (81)
To compare the last two displays (80) and (81), we may write
$\binom{k^{\prime}+k^{\prime\prime}}{k^{\prime}}b_{n-1,k^{\prime}+k^{\prime\prime}+1}=\sum_{j\geq 0}\binom{n-2}{j}b_{j+1,k^{\prime}+1}b_{n-2-j,k^{\prime\prime}}\\ =\frac{R^{\prime}}{R}\delta_{k^{\prime},0}b_{n-2,k^{\prime\prime}}+\sum_{j\geq 1}b_{j+1,k^{\prime}+1}b_{n-2,k^{\prime\prime}+1}^{[j]}$ (82)
upon interpreting the lhs as a sum over partitions of $\{1,\ldots,n-1\}$ into $k^{\prime}+k^{\prime\prime}+1$ blocks, $k^{\prime}+1$ of which are marked, the element $n-1$ being always in a marked block: the first equality is obtained by summing over the number $j$ of elements other than $n-1$ that are in a marked block, and the second equality is obtained by noting that $\binom{n-2}{j}b_{n-2-j,k^{\prime\prime}}=b_{n-2,k^{\prime\prime}+1}^{[j]}$ for $j\geq 1$. We deduce that the sum appearing at the end of (81), with $k$ replaced by $k^{\prime\prime}$, can be rewritten as
$\sum_{j\geq 1}\sum_{k^{\prime}\geq 0}k^{\prime}!q_{k^{\prime}}(\ell_{n+1})b_{j+1,k^{\prime}+1}b_{n-2,k^{\prime\prime}+1}^{[j]}=\sum_{k^{\prime}\geq 0}\binom{k^{\prime}+k^{\prime\prime}}{k^{\prime}}k^{\prime}!q_{k^{\prime}}(\ell_{n+1})b_{n-1,k^{\prime}+k^{\prime\prime}+1}\\ -\frac{R^{\prime}}{R}b_{n-2,k^{\prime\prime}}$ (83)
and hence
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}-D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1}-1}R^{\prime}\times\\ \sum_{k\geq 0}k!p_{k}(\ell_{1},\ldots,\ell_{n})\left(b_{n-2,k}-(\ell_{1}+\cdots+\ell_{n}-k-1)b_{n-2,k+1}\right).$ (84)
By comparing with (75) we see that
$\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}-D_{2\ell_{n+1}}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=R^{\ell_{n+1}}\left(\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n},0}\right|_{\mathrm{bip}}-\frac{\partial}{\partial t}\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\right)$ (85)
and we immediately deduce that (78) is again a direct consequence of the string equation (77).
Remark 3.8.
We may similarly check that the explicit expression (44) for $\left.T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}$ in the planar quasi-bipartite case is also consistent with the relations (59) and (68) for the addition of an extra boundary. This property is now a direct consequence of the string equation relating $\tilde{p}_{k+1}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$ to $\tilde{p}_{k}(\ell_{1},\ell_{2};\ell_{3},\ldots,\ell_{n})$, see [polytightmaps, Proposition 2.10].
4 The quasi-polynomiality phenomenon
We have seen in Section 2.4 that the generating function of planar bipartite maps with (at least three) tight boundaries is an even polynomial in the boundary lengths times a power of the basic generating function $R$. As mentioned in the introduction, this is a particular case of a more general quasi-polynomiality phenomenon.
Precisely, given a positive integer $n$, let us call parity-dependent quasi-polynomial in $n$ variables a function $f$ of $n$ integer variables, such that there exists a (necessarily unique) family of polynomials $(f_{I})_{I}$ indexed by subsets $I$ of $\{1,\ldots,n\}$, such that $f(m_{1},\ldots,m_{n})=f_{I}(m_{1},\ldots,m_{n})$ whenever the $m_{i}$ with $i\in I$ are odd integers, and the $m_{i}$ with $i\notin I$ are even integers. By coefficients of $f$, we mean the collection of coefficients of the polynomials $f_{I}$, which may belong to an arbitrary ring; by total degree of $f$, we mean the maximum of the total degrees of the $f_{I}$. It is convenient to allow the value $n=0$, in which case it is understood that $f$ is just a constant element of the ring at hand.
In order to state the main result of this section, we also introduce the quantity
$\chi(\mathcal{M}_{g,n}):=\begin{cases}(-1)^{n-1}(n-3)!&\text{if $g=0$,}\\ (-1)^{n-1}\frac{(2g+n-3)!}{(2g-2)!}\zeta(1-2g)&\text{if $g>0$,}\end{cases}$ (86)
which, for $02-2g-n<$ 0, is the Euler characteristic of the moduli space $\mathcal{M}_{g,n}$ [Harer1986].
Theorem 4.1.
Let $g,n$ be nonnegative integers such that $02-2g-n<$ 0. Then, there exists a parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, of total degree $3g-3+n$, such that for any nonnegative integers $\ell_{1},\ldots,\ell_{n}$, we have
$T^{(g)}_{\ell_{1},\ldots,\ell_{n}}=R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})-\chi(\mathcal{M}_{g,n})t^{2-2g-n}\delta_{\ell_{1}+\cdots+\ell_{n},0}\,.$ (87)
The coefficients of $\mathfrak{T}^{(g)}_{n}$ are rational functions with rational coefficients of the quantities $\frac{R^{(k)}}{R}$ and $\frac{S^{(k)}}{\sqrt{R}}$, $k=1,\ldots,3g-2+n$, where $R^{(k)},S^{(k)}$ denote the $k$-th derivatives of the series $R,S$ with respect to the vertex weight $t$.
The quasi-polynomiality phenomenon was first observed in the language of topological recursion by Norbury and Scott [NoSc13], who showed that the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the series expansion (10) of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ are quasi-polynomials in $\ell_{1},\ldots,\ell_{n}$ which are odd in each variable. We review their result in Section 4.1: combining it with the trumpet decomposition yields (87) in the case where $n$ and all $\ell_{i}$ are non-zero.
However, the quasi-polynomiality of $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ does not imply a priori that of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, as there may be pathologies when some variables are set to 00. To circumvent this problem, and show that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ indeed coincides with an even quasi-polynomial except when all variables are set to 00, we will make use in Section 4.2 of the tight boundary insertion relations established in Section 3. The proof is by induction on the number of boundaries $n$, and we still use results from topological recursion to initialise the induction.
Remark 4.2.
For $n=0$, Theorem 4.1 states that for any $g\geq 2$, the generating function $F^{(g)}$ of maps of genus $g$ without boundaries reads
$F^{(g)}=\mathfrak{F}^{(g)}-\chi(\mathcal{M}_{g,0})t^{2-2g}\,,$ (88)
with $\mathfrak{F}^{(g)}$ a rational function with rational coefficients of $\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}}$, $k=1,\ldots,3g-2$. As we will see below, this result is obtained by combining the property known from topological recursion that $F^{(g)}$ can be expressed in terms of the so-called moments, with an expression of the moments in terms of the derivatives of $R$ and $S$.
Remark 4.3.
When setting the weights $t_{1},t_{2},\ldots$ for inner faces to zero, i.e. when we consider tight maps for which every face is a tight boundary, then $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ reduces to $t^{2-2g-n}$ times Norbury’s lattice count (quasi-)polynomial $N_{g,n}(\ell_{1},\ldots,\ell_{n})$ [Norbury2010]. It is shown in this reference that $N_{g,n}(0,\ldots,0)=\chi(\mathcal{M}_{g,n})$. This is consistent, in view of (87), with the fact that $T^{(g)}_{0,\ldots,0}$ vanishes when we set the weights for inner faces to zero (there are no tight maps having only boundary-vertices and no faces).
4.1 Maps without boundary-vertices
In this section, we establish the quasi-polynomiality property (87) in the case where $n$ and all $\ell_{i}$ are non-zero. We will do so by combining Theorem 2.5 with known results coming from the theory of topological recursion. One reason behind the fact that this theory is formulated in terms of $\omega^{(g)}_{n}$ instead of $W^{(g)}_{n}$ is that the former are rational functions of their $n$ variables with a simple pole structure, see for instance [Eynard2016, Section 3.3.1]. It was recognized by Norbury and Scott that this simple pole structure amounts to the following proposition:
Proposition 4.4.
For $g\geq 0$, $n\geq 1$ and $02-2g-n<$ 0, the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the expansion (10) of $\omega^{(g)}_{n}$ are of the form
$\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\ell_{1}\cdots\ell_{n}\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ (89)
with $\mathfrak{T}^{(g)}_{n}$ a parity-dependent quasi-polynomial in $n$ variables of total degree $3g-3+n$.
The above proposition is a reformulation of [NoSc13, Theorem 1] [^5] Precisely, this theorem considers the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ around $z_{i}=0$. Here, our $\hat{\tau}_{\ell_{1},\ldots,\ell_{n}}^{(g)}$ are rather defined in terms of the expansion around $z_{i}=\infty$. But the two expansions can be related through antisymmetry properties of $\omega^{(g)}_{n}$, see (94)., in which the quantity $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ is denoted by $N^{g}_{n}(\ell_{1},\ldots,\ell_{n})$. Note that we do not characterize the structure of the coefficients of $\mathfrak{T}^{(g)}_{n}$ at this stage. The relation (89) implies that (87) holds when $n$ and all $\ell_{i}$ are non-zero, by (11) and (17).
For completeness, let us sketch a proof of Proposition 4.4 based on the results available in [Eynard2016]. We start with a technical lemma.
Lemma 4.5.
Let $P(n)$ be a polynomial of degree at most $12d+$ 1, which is equivalent to the series $F(z):=\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}$ being a rational function of the form $N(z)/(z-1)^{2d+2}$, with $N$ a polynomial of degree at most $12d+$ 1. Then, the following properties are equivalent:
  • (i)
    $P$ is odd,
  • (ii)
    $F$ is such that $F(z)=z^{-2}F(z^{-1})$,
  • (iii)
    $N$ satisfies $N(z)=z^{2d}N(z^{-1})$, i.e. it is a self-reciprocal (palindromic) polynomial.
Proof.
By [StanleyEC1, Proposition 4.2.3] [^6] Short rederivation: we consider the vector space $\mathcal{L}$ of formal Laurent series of the form $\sum_{n\in\mathbb{Z}}a_{n}z^{n}$. We cannot multiply two elements of this space in general, but we can multiply an element of $\mathcal{L}$ by a polynomial in $z$, and this operation is linear. Let us consider in particular the series $\sum_{n\in\mathbb{Z}}\frac{P(n)}{z^{n+1}}\in\mathcal{L}$, and the polynomial $(z-1)^{2d+2}$. Their product vanishes since $P$ is a polynomial of degree at most $12d+$ 1 hence is annihilated by the $(2d+2)$-th power of the forward difference operator. We deduce that $(z-1)^{2d+2}\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}=-(z-1)^{2d+2}\sum_{n\leq-1}\frac{P(n)}{z^{n+1}}$. The left-hand side may be interpreted as a product in the ring $\mathbb{C}((z^{-1}))$ and is thus equal to $N(z)\in\mathbb{C}[z]$. Therefore the right-hand side is also equal to $N(z)$, and we obtain an identity valid in $\mathbb{C}[[z]]$. Multiplying by the inverse of $(z-1)^{2d+2}$ in $\mathbb{C}[[z]]$ (which differs from that in $\mathbb{C}((z^{-1}))$!) and changing $n$ in $-n$, we obtain the wanted series identity., the rational function $F(z)$ admits around 00 the expansion $F(z)=\allowbreak-\sum_{n\geq 1}P(-n)z^{n-1}$ hence we have
$z^{-2}F(z^{-1})=-\sum_{n\geq 1}\frac{P(-n)}{z^{n+1}}.$ (90)
This implies the equivalence between (i) and (ii). The equivalence between (ii) and (iii) is straightforward. ∎
Proof of Proposition 4.4.
According to [Eynard2016, Lemma 3.3.3 page 89, Theorem 3.5.1 page 121, and last equation on page 140] [^7] There seems to be a discrepancy between Theorem 3.5.1 and the equation on page 140: $\sum d_{i}$ is at most $6g-6+2n$ in the former, and $6g-4+2n$ in the latter. We assume that the former is correct since it holds for pairs of pants ($g=0,n=3$) and lids ($g=1,n=1$)., $\omega^{(g)}_{n}$ has a partial fraction decomposition of the form
$\omega^{(g)}_{n}(z_{1},\ldots,z_{n})=\sum_{\epsilon\in\{+1,-1\}^{n}}\sum_{\begin{subarray}{c}d\in\mathbb{N}^{n}\\ d_{1}+\cdots+d_{n}\leq 6g-6+2n\end{subarray}}\frac{c_{\epsilon,d}}{(z_{1}-\epsilon_{1})^{d_{1}+2}\cdots(z_{n}-\epsilon_{n})^{d_{n}+2}}$ (91)
where the $c_{\epsilon,d}$ are some coefficients depending on the spectral curve. This implies that, in the expansion (10) of $\omega^{(g)}_{n}$, the $\hat{\tau}$ coefficients are of the form
$\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\sum_{\epsilon\in\{+1,-1\}^{n}}Q^{(g,\epsilon)}(\ell_{1},\ldots,\ell_{n})\epsilon_{1}^{\ell_{1}}\cdots\epsilon_{n}^{\ell_{n}}.$ (92)
where $Q^{(g,\epsilon)}$ is a polynomial of degree at most $6g-6+3n$. It remains to check that these polynomials are indeed odd in each variable, so that we may write, for some polynomials $P^{(g,\epsilon)}$ of degree at most $3g-3+n$,
$Q^{(g,\epsilon)}(\ell_{1},\ldots,\ell_{n})=\ell_{1}\cdots\ell_{n}\,P^{(g,\epsilon)}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\,.$ (93)
For this we use the antisymmetry property of [Eynard2016, Lemma 3.3.2, page 88], which entails that
$\omega^{(g)}_{n}(z_{1},z_{2},\ldots,z_{n})=z_{1}^{-2}\omega^{(g)}_{n}(z_{1}^{-1},z_{2},\ldots,z_{n})$ (94)
and similarly for all other variables. In the expansion (91), each term of the sum over $\epsilon$ must be separately invariant under the antisymmetry mapping $f(z_{i})\mapsto z_{i}^{-2}f(z_{i}^{-1})$. Such term corresponds to the term $Q^{(g,\epsilon)}(\ell_{1},\ldots,\ell_{n})\epsilon_{1}^{\ell_{1}}\cdots\epsilon_{2}^{\ell_{2}}$ in (92) and we claim that $Q^{(g,\epsilon)}$ is odd in each of its $n$ variables. Indeed, by the transformation $z_{i}\mapsto\epsilon_{i}z_{i}$, it suffices to consider the case $\epsilon=(1,\ldots,1)$, and the claim results from Lemma 4.5 and induction over $n$. As $m_{i}=\ell_{i}^{2}$ has the same parity as $\ell_{i}$, we conclude that (89) holds with
$\mathfrak{T}^{(g)}_{n}(m_{1},\ldots,m_{n})=\sum_{\epsilon\in\{+1,-1\}^{n}}P^{(g,\epsilon)}(m_{1},\ldots,m_{2})\epsilon_{1}^{m_{1}}\cdots\epsilon_{n}^{m_{n}}\,,$ (95)
which is indeed a parity-dependent quasi-polynomial in $n$ variables of total degree $3g-3+n$. ∎
4.2 Maps with boundary-vertices
The purpose of this section is to complete the proof of Theorem 4.1, by including the case where some of the boundary lengths $\ell_{1},\ldots,\ell_{n}$ vanish, and by showing the structure properties of the coefficients of the quasi-polynomial $\mathfrak{T}^{(g)}_{n}$. We proceed by induction on $n$, by using the recursion relations of Section 3 for the induction step. As a warmup, let us first consider the easier case of maps with even face degrees. Recall the relation (23) between the $T$’s and the $\tau$’s, and the notation $\left.\right|_{\mathrm{bip}}$ from Section 2.4 which indicates that we set the weights $t_{1},t_{3},\dots$ for inner faces of odd degrees to zero.
Proposition 4.6.
Let $g,n$ be nonnnegative integers such that $02-2g-n<$ 0. There exists an element $Q^{(g)}_{n}(x_{1},\ldots,x_{3g-2+n};u_{1},\ldots,u_{n})\in\mathbb{Q}(x_{1},\ldots,x_{3g-2+n})[u_{1},\ldots,u_{n}]$ and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have
$\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}=Q^{(g)}_{n}\left(\frac{R^{\prime}}{R},\ldots,\frac{R^{(3g-2+n)}}{R};\ell_{1}^{2},\ldots,\ell_{n}^{2}\right)-c(g,n)t^{2-2g-n}\delta_{\ell_{1}+\cdots+\ell_{n},0}\,.$ (96)
Proof.
As mentioned above, we proceed by induction on $n$ for a fixed value of $g$. The induction step relies on the recursion relations of Propositions 3.3 and 3.5 which, for $t_{1}=t_{3}=\cdots=0$ (so that $S=0$), reduce to
$\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\,.$ (97)
Here, $D_{m}$ is the tight boundary insertion operator defined at (61), reading for $S=0$
$D_{2\ell}=\begin{cases}\sum_{i=1}^{\ell}(-1)^{\ell+i}\binom{\ell+i-1}{\ell-i}R^{\ell-i}\frac{\partial}{\partial t_{2i}}&\text{if $\ell>0$,}\\ \frac{\partial}{\partial t}&\text{if $\ell=0$.}\end{cases}$ (98)
In the second line of (97), it is understood that, if $\ell_{i}=0$, then the corresponding sum over $0<m_{i}<\ell_{i}$ vanishes. In terms of the $\tau$’s, this relation rewrites as
$\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=\frac{D_{2\ell_{n+1}}\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}}{R^{\ell_{n+1}}}\\ +\frac{R^{\prime}}{R}\sum_{i=1}^{n}\left(\ell_{i}\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\sum_{0<m_{i}<\ell_{i}}(2m_{i})\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\right)\,.$ (99)
Here, we have used the Leibniz product rule for the differential operator $D_{2\ell_{n+1}}$, and the relation $D_{2\ell}R=R^{\ell}R^{\prime}$ which is the case $j=0$ of (79), recalling the definition (33) of $b_{n,k}$.
Let us now sketch the rough idea of the induction, ignoring the pathological term of (96) for now:
  • from the recursion relation (99), it is relatively straightforward that if $\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}$ is polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$, then $\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}$ is also polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$: it is clear for the first term since the linear operator $D_{2\ell_{n+1}}$ does not act on these variables, while the second term involves a summation which turns out to preserve the wanted polynomiality property,
  • regarding the new variable $\ell_{n+1}$, we use the assumption that the coefficients of $\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}$ are rational in the derivatives of $R$: acting on this quantity with the differential operator $D_{2\ell_{n+1}}$ and using the chain rule, it follows from (79) that $R^{-\ell_{n+1}}D_{2\ell_{n+1}}\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}$ is polynomial in $\ell_{n+1}^{2}$, also with rational coefficients in the derivatives of $R$.
Let us now make this argument precise. Assuming that the expression (96) holds at rank $n$, we plug it into the right-hand side of (99), and analyse the two terms separately. The second term is easier to deal with, since it does not depend on $\ell_{n+1}$ nor receives a contribution from the pathological term of (96). It equals
$\frac{R^{\prime}}{R}\sum_{i=1}^{n}\left(\ell_{i}Q^{(g)}_{n}\left(\ldots,\ell_{i}^{2},\ldots\right)+\sum_{0<m_{i}<\ell_{i}}(2m_{i})Q^{(g)}_{n}\left(\ldots,m_{i}^{2},\ldots\right)\right)\,,$ (100)
where, for readability, we only display the $u_{i}$ variable of $Q^{(g)}_{n}$. Now, it is an exercise (solved in Lemma B.1 below) that for every univariate polynomial $\mathfrak{q}$, the quantity $\ell\mathfrak{q}(\ell^{2})+\sum_{0<m<\ell}(2m)\mathfrak{q}(m^{2})$ is again a polynomial in $\ell^{2}$. Therefore, the second term of (99) is of the wanted form. To deal with the first term, we apply the chain rule to obtain
$\frac{D_{2\ell_{n+1}}\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}}{R^{\ell_{n+1}}}=\sum_{j=1}^{3g-2+n}\frac{D_{2\ell_{n+1}}\left(\frac{R^{(j)}}{R}\right)}{R^{\ell_{n+1}}}\frac{\partial Q^{(g)}_{n}\left(\ldots\right)}{\partial x_{j}}-\\ c(g,n)\frac{D_{2\ell_{n+1}}t^{2-2g+n}}{R^{\ell_{n+1}}}\delta_{\ell_{1}+\cdots+\ell_{n},0}\,.$ (101)
Here, the first term is clearly polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, it is also a polynomial in $\ell_{n+1}^{2}$ by (79), which also entails that the coefficients are rational in $R^{(k)}/R$ for $1\leq k\leq 3g-2+(n+1)$. Finally, the second term in the last display is zero unless $\ell_{1},\ldots,\ell_{n+1}$ all vanish, in which case it is equal to $-c(g,n)\frac{\partial}{\partial t}t^{2-2g-n}=-c(g,n+1)t^{2-2g-(n+1)}$, where $c(g,n+1)=(2-2g+n)c(g,n)$. This completes the proof of the induction step: if the statement of Proposition 4.6 is true for $(g,n)$, then it is also true for $(g,n+1)$.
It remains to initialise the induction. For $g=0$, we initialise the induction at $n=3$ by using the explicit formula of [triskell, Theorem 1.1], which is the bipartite version of Theorem A.1 in Appendix A. For $g\geq 1$, we rely on consequences of the topological recursion that are described in [budd2020irreducible]. Assume first that $g\geq 2$, in which case we initialise the induction at $n=0$: in view of Remark 4.2, we only have to check that the generating function $\left.F^{(g)}\right|_{\mathrm{bip}}=\left.\tau^{(g)}_{\varnothing}\right|_{\mathrm{bip}}$ of essentially bipartite maps of genus $g$ without boundary is a rational function of $R^{(k)}/R,1\leq k\leq 3g-2$. By [budd2020irreducible, Equations (31) and (32)], there exists a polynomial $\tilde{P}_{g}$ with rational coefficients in $33g-$ 3 variables such that for $t=1$ we have
$\left.F^{(g)}\right|_{\mathrm{bip}}[t=1]=\tilde{P}_{g}(0,\ldots,0)-\left(\frac{1}{\bar{M}_{0}}\right)^{2g-2}\tilde{P}_{g}\left(\frac{\bar{M}_{1}}{\bar{M}_{0}},\ldots,\frac{\bar{M}_{3g-3}}{\bar{M}_{0}}\right)\,,$ (102)
where $\bar{M}_{p}$ are the so-called moments, defined at [budd2020irreducible, Equation (30)]. Using Euler’s relation, we can restore the dependence on $t$ by the formula
$\left.F^{(g)}\right|_{\mathrm{bip}}=t^{2-2g}\left.F^{(g)}\right|_{\mathrm{bip}}[t=1](t_{2},tt_{4},t^{2}t_{6},\ldots)\,.$ (103)
Consequently, letting $\bar{M}_{p}^{(0)}=t\bar{M}_{p}(t_{2},tt_{4},t^{2}t_{6},\ldots)$, in accordance with [budd2020irreducible, Equation (43)], we have
$\left.F^{(g)}\right|_{\mathrm{bip}}=t^{2-2g}\tilde{P}_{g}(0,\ldots,0)-\left(\frac{1}{\bar{M}_{0}^{(0)}}\right)^{2g-2}\tilde{P}_{g}\left(\frac{\bar{M}_{1}^{(0)}}{\bar{M}_{0}^{(0)}},\ldots,\frac{\bar{M}_{3g-3}^{(0)}}{\bar{M}_{0}^{(0)}}\right)\,.$ (104)
Finally, [budd2020irreducible, Lemma 9] shows that $\bar{M}_{0}^{(0)}=R/R^{\prime}$, while we have
$\frac{\bar{M}_{p}^{(0)}}{\bar{M}_{0}^{(0)}}=\frac{\mathsf{T}_{p}\left(\frac{R^{\prime}}{R},\ldots,\frac{R^{(p+1)}}{R}\right)}{(R^{\prime}/R)^{2p}}$ (105)
for some homogeneous polynomial $\mathsf{T}_{p}$ of degree $2p$. This entails that $\left.F^{(g)}\right|_{\mathrm{bip}}$ is of the wanted form (96): the second term in (104) is of the form $Q^{(g)}_{0}\left(\frac{R^{\prime}}{R},\ldots,\frac{R^{(3g-2)}}{R}\right)$ with $Q^{(g)}_{0}\in\mathbb{Q}(x_{1},\ldots,x_{3g-2})$ (there are no variables $u_{i}$ for $n=0$), and the first term is of the form $t^{2-2g}c(g,0)$ with $c(g,0)=-\tilde{P}_{g}(0,\ldots,0)$.
In the case $g=1$, we rely on [budd2020irreducible, Equation (31)], stating, in our notation, that $F^{(1)}=\ln(tR^{\prime}/R)/12$. This entails, again by the boundary insertion formula and by (79), that
$\left.\tau^{(1)}_{2\ell}\right|_{\mathrm{bip}}=\frac{D_{2\ell}F^{(1)}}{R^{\ell}}=\frac{D_{2\ell}\left(\frac{R^{\prime}}{R}\right)}{12\frac{R^{\prime}}{R}R^{\ell}}+\frac{\delta_{\ell,0}}{12t}\,=\frac{1}{12}\left((\ell^{2}-1)\frac{R^{\prime}}{R}+\frac{R^{\prime\prime}}{R^{\prime}}\right)+\frac{\delta_{\ell,0}}{12t}$ (106)
which is of the wanted form (96), with $c(1,1)=-1/12$. This concludes the proof of Proposition 4.6. ∎
Let us record in passing the following interesting consequence of (106), which calls for a bijective interpretation.
Proposition 4.7.
The generating function of essentially bipartite maps of genus $1$ with one tight boundary of length $2\ell$ is given by
$\left.T^{(1)}_{2\ell}\right|_{\mathrm{bip}}=\frac{R^{\ell}}{12}\left((\ell^{2}-1)\frac{R^{\prime}}{R}+\frac{R^{\prime\prime}}{R^{\prime}}\right)+\frac{\delta_{\ell,0}}{12t}\,.$ (107)
We now state the non bipartite analogue of Proposition 4.6.
Proposition 4.8.
Let $g,n$ be nonnnegative integers such that $02-2g-n<$ 0. There exists a parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, of total degree $3g-3+n$, whose coefficients are rational functions with rational coefficients of the quantities $\frac{R^{(k)}}{R}$ and $\frac{S^{(k)}}{\sqrt{R}}$ for $k=1,\ldots,3g-2+n$, and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have
$\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}=\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})-c(g,n)t^{2-2g-n}\delta_{\ell_{1}+\cdots+\ell_{n},0}\,.$ (108)
The proof, by induction, is in the same spirit as that of Proposition 4.6, but involves extra difficulties. For this reason, we offload it to the appendices: Appendix B is devoted to the induction step and Appendix C to the initialisation.
End of the proof of Theorem 4.1.
In view of Proposition 4.8 and of the relation (23) between the $T$’s and the $\tau$’s, the only task left is to show that $c(g,n)=\chi(\mathcal{M}_{g,n})$. But, as discussed in Remark 4.3, this is a consequence of [Norbury2010, Theorem 2] which asserts that the lattice count polynomial $N_{g,n}$ evaluates to $\chi(\mathcal{M}_{g,n})$ when all its variables are set to zero. ∎
5 Conclusion and discussion
In this paper, we have seen that the generating function $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ of maps of genus $g$ with $n$ tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$ appears within the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$, a fundamental quantity of the theory of topological recursion. This coincidence is combinatorially explained by the trumpet decomposition. We have found recursion relations, also of combinatorial nature, expressing $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$ in terms of $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Finally, we have seen that, when at least one $\ell_{i}$ is non-zero, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ is equal to a parity-dependent quasi-polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ times a simple power of $R$.
At this stage, a tantalizing question is whether one can derive bijectively an expression for $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. So far, only the situation $(g,n)=(0,3)$ has been treated in [triskell] for the essentially bipartite case, and in Appendix A for the general case. In particular, the remarkably simple formulas (34) for $(g,n)=(0,4)$ and (107) for $(g,n)=(1,1)$, in the case of maps with faces of even degree only, are still waiting for a bijective proof. One may also attempt to find a combinatorial interpretation of (32), given that its constituents $p_{k}(\ell_{1},\ldots,\ell_{n})$ and $b_{n-2,k+1}$ have themselves interpretations in terms of tight maps [polytightmaps] and set partitions, respectively.
Acknowledgements.
This work is supported by the Agence Nationale de la Recherche via the grants ANR-18-CE40-0033 “Dimers”, ANR-19-CE48-0011 “Combiné” and ANR-23-CE48-0018 “CartesEtPlus”. JB acknowledges the hospitality of the Laboratoire de Physique of ENS de Lyon during the initial stages of this work.
Appendix A Bijective enumeration of tight pairs of pants
The purpose of this appendix is to establish the following:
Theorem A.1.
Let $a$, $b$ and $c$ be nonnegative integers or half-integers. Then, the generating function $T^{(0)}_{2a,2b,2c}$ of planar maps with three labeled distinct tight boundaries of lengths $2a$, $2b$, $2c$, counted with a weight $t$ per vertex different from a boundary-vertex and, for all $k\geq 1$, a weight $t_{k}$ per inner face of degree $k$, is given by
$T^{(0)}_{2a,2b,2c}=\begin{cases}R^{-1}\frac{\partial R}{\partial t}-t^{-1}&\text{if $a=b=c=0$,}\\ R^{a+b+c-1}\frac{\partial R}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}$,}\\ R^{a+b+c-\frac{1}{2}}\frac{\partial S}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}-\frac{1}{2}$,}\end{cases}$ (109)
where $R$ and $S$ are the formal power series in $t,t_{1},t_{2},t_{3},\ldots$ defined in Section 1.2.
This theorem is an extension of [triskell, Theorem 1.1] to non bipartite maps. Note that, in this reference (which the reader is invited to consult for any definition not found in the current paper), we used the notation $T_{a,b,c}$ instead of the present notation $T^{(0)}_{2a,2b,2c}$. The (essentially) bipartite case is recovered upon taking $t_{1}=t_{3}=t_{5}=\cdots=0$ which, as noted above, implies that $R$ satisfies (4) and that $S$ vanishes. In that case, $T^{(0)}_{2a,2b,2c}$ is non-zero if and only if $a+b+c$ is an integer, as wanted.
To prove Theorem A.1, we will use the well-known fact that any map $M$ can be transformed into a bipartite map by doing a subdivision of every edge, that is adding a new vertex at the midpoint of every edge, which is thus split into two. We denote by $\phi(M)$ the resulting subdivided map. The vertices of $\phi(M)$ consist of regular vertices (those originally in $M$) and midpoint vertices (those added by the subdivision, which have degree $2$). The faces of $\phi(M)$ are those of $M$, only their degree is doubled.
Figure 5: A summary representation of the bijection of [triskell] between maps with three tight boundaries and quintuples $(D_{1},D_{2},D_{3},T_{1},T_{2})$ made of three bigeodesic diangles $D_{1},D_{2},D_{3}$ and two bigeodesic triangles $T_{1},T_{2}$. The solid lines indicate the identification between attachment points for type I. For type II, the two outer identification lines have to be replaced by the dotted lines.
We then proceed by “conjugating” the bijective decomposition of [triskell] by $\phi$. Indeed, let $M$ be a planar map with three tight boundaries of lengths $2a,2b,2c$: clearly, $\phi(M)$ is a bipartite planar map with three tight boundaries. We may then apply one of two possible decompositions described in [ibid., Sections 5.6 and 5.7], namely:
  • the decomposition of type I if $a\leq b+c$ and $b\leq c+a$ and $c\leq a+b$,
  • the decomposition of type II if $a\geq b+c$ or $b\geq c+a$ or $c\geq a+b$.
In both cases, we obtain a quintuple consisting of five pieces: two so-called bigeodesic triangles, and three bigeodesic diangles. The vertices of these pieces are those of $\phi(M)$, and we see that a midpoint vertex of $\phi(M)$ remains of degree $2$ in each piece to which it belongs: indeed, the decomposition is made by cutting along bigeodesics, and distinct bigeodesics may only merge at vertices of degree at least $3$, which are necessarily regular. As a consequence of this remark, each piece admits a (unique) preimage by $\phi$, and so our decomposition amounts to cutting $M$ itself into five pieces, denoted $(D_{1},D_{2},D_{3},T_{1},T_{2})$ on Figure 5.
While this approach seems very natural, the main difficulty is to characterize precisely which sorts of pieces we may obtain in the decomposition. Recall from [ibid., Sections 2.3 and 2.4] that bigeodesic diangles and triangles are planar maps with one boundary-face having several (four and six, respectively) distinguished incident corners, half of them being the so-called attachment points. Let us here call tips the other distinguished corners: by the above remark, we see that the tips of the pieces obtained in the decomposition of $\phi(M)$ are always incident to regular vertices. However, the attachment points may be incident to either regular or midpoint vertices. We are thus led to introduce the following definitions.
In a map $M$, a midpoint corner is a corner of the subdivided map $\phi(M)$ which is incident to a midpoint vertex. A regular corner is a corner in the usual sense, it corresponds to a corner of $\phi(M)$ incident to a regular vertex. We define the distance $d_{M}(c,c^{\prime})$ between two (regular or midpoint) corners $c,c^{\prime}$ of $M$ as half the graph distance between their incident vertices in $\phi(M)$. The distance between two regular corners, or between two midpoint corners, is thus an integer, while the distance between a regular corner and a midpoint corner is a half-integer. The notion of geodesic boundary interval of [ibid., Section 2.1] extends naturally to midpoint corners.
Figure 6: Schematic representation of a generalized bigeodesic diangle (left) and triangle (right). The tips $c_{1},c_{2},c_{3}$ are regular corners, incident to vertices (shown as solid disks). The attachment points $c_{12},...$ are either regular corners or midpoint corners, incident to midpoint vertices of the subdivided map (shown as small squares). We use the same red/blue coloring conventions as [triskell, Figures 2.2 and 2.5].
A generalized bigeodesic diangle $D$ is defined as in [ibid., Section 2.3] except that we no longer require $D$ to be bipartite, and that we allow that zero, one or two among the attachment points $c_{12},c_{21}$ are midpoint corners. We still require that the tips $c_{1},c_{2}$ are regular corners of $D$. See the left of Figure 6 for an illustration. The exceedance of $D$ is defined as
$e(D):=d_{D}(c_{12},c_{1})-d_{D}(c_{21},c_{1})=d_{D}(c_{21},c_{2})-d_{D}(c_{12},c_{2}).$ (110)
This quantity is half the exceedance of $\phi(D)$, which is clearly a bigeodesic diangle in the original sense. When $e(D)$ is an integer, the two attachment points are of the same kind (regular or midpoint), and they are of different kinds when $e(D)$ is a half-integer.
A generalized bigeodesic triangle $T$ is defined as in [ibid., Section 2.4] except that we no longer require $T$ to be bipartite, and that we allow the attachment points $c_{12},c_{23},c_{31}$ to be midpoint corners. We still require that the tips $c_{1},c_{2},c_{3}$ are regular corners of $T$. Recall that the definition of a bigeodesic triangle implies that $d_{T}(c_{12},c_{1})=d_{T}(c_{31},c_{1})$ (and circular permutations thereof), and hence all the attachment points must be of the same kind. The triangle is said odd if $c_{12},c_{23},c_{31}$ are midpoint corners, and even if they are regular corners. See the right of Figure 6 for an illustration. Clearly, $\phi(T)$ is a bigeodesic triangle in the usual sense.
With these new definitions at hand, by applying [ibid., Theorem 3.1] to $\phi(M)$, we arrive at the following:
Proposition A.2.
Let $M$ be a planar map with three tight boundaries of lengths $2a,2b,2c$.
If $a\leq b+c$, $b\leq c+a$, and $c\leq a+b$, then by the type I decomposition of [triskell, Section 5.6], $M$ is in one-to-one correspondence with a quintuple $(D_{1},D_{2},D_{3},T_{1},T_{2})$ where $D_{1},D_{2},D_{3}$ are generalized bigeodesic triangles of respective exceedances $e(D_{1})=a+c-b$, $e(D_{2})=a+b-c$, $e(D_{3})=b+c-a$, and $T_{1},T_{2}$ are generalized bigeodesic triangles, with the constraints that the attachment points which are identified together on Figure 5 must be of the same (regular or midpoint) kind, and that not all elements of this quintuple are reduced to the vertex-map.
If $b\geq a+c$, say, then the type II decomposition of [ibid., Section 5.7] produces a quintuple $(D_{1},D_{2},D_{3},T_{1},T_{2})$ with the same properties, except that the exceedances are now $e(D_{1})=b-a-c$, $e(D_{2})=2a$, $e(D_{3})=2c$.
For both type I and type II, we have the relation
$e(D_{1})+e(D_{2})+e(D_{3})=a+b+c\,.$ (111)
The above proposition encompasses a number of different cases. When $a+b+c$ is an integer, all the exceedances are integers, and therefore all twelve attachment points are of the same kind. If they are all regular (resp. midpoint) corners, then $T_{1}$ and $T_{2}$ are both even (resp. odd) triangles. When $a+b+c$ is a half-integer, $e(D_{1})$ is a half-integer and hence $T_{1}$ and $T_{2}$ have different “parities”. The other exceedances $e(D_{2})$ and $e(D_{3})$ are half-integers for type I, and integers for type II.
Let us now discuss the enumerative consequences of Proposition A.2. To this end, it is necessary to introduce the generating functions of the various objects that may arise in the situations described above. Generally speaking, we assign a weight $t_{k}$ per inner face of degree $k$ (before subdivision), and a weight $t$ per vertex (of the map, i.e. per regular vertex of the subdivided map). Following the conventions of [ibid.], vertices that lie on strictly geodesic boundaries (displayed in red on the figures), and that are not incident to attachment points, do not receive the weight $t$. We denote by $Y$ (resp. $\widetilde{Y}$) the generating function of generalized bigeodesic triangles that are even (resp. odd). We then treat the case of diangles in the following proposition.
Proposition A.3.
Let $X$ denote the generating function of generalized bigeodesic diangles with exceedance 00, whose attachment points are both regular corners. Then the following holds.
The generating function of generalized bigeodesic diangles with exceedance $e\in\mathbb{Z}_{\geq 0}$, whose attachment points are both regular corners, is equal to $R^{e}X$.
The generating function of generalized bigeodesic diangles with exceedance $e\in\mathbb{Z}_{\geq 0}+\frac{1}{2}$, where the first attachment point is a regular corner and the second a midpoint corner, is equal to $R^{e+\frac{1}{2}}X/t$.
The generating function of generalized bigeodesic diangles with exceedance $e\in\mathbb{Z}_{\geq 0}$, whose attachment points are both midpoint corners, is equal to $R^{e+1}X/t^{2}$.
Proof.
The first statement is obtained by a straightforward adaptation of the proof of [ibid., Proposition 2.2], noting that the arguments do not require that the maps be bipartite, and that $R$ can still be interpreted as the generating function for elementary slices.
The second statement follows from the first one, by observing that there is a bijection between the generalized bigeodesic diangles whose exceedance is $e\in\mathbb{Z}_{\geq 0}+\frac{1}{2}$ and whose second attachment point is a midpoint corner, and diangles whose exceedance is $e+\frac{1}{2}$ and whose both attachment points are regular corners. Indeed, looking at Figure 6, we simply have to move the second attachment point $c_{21}$ to the nearest (red) regular vertex in the direction of $c_{1}$, and we need to divide by $t$ to unweigh this regular vertex.
Finally, the third statement follows from a similar observation: we now move the two attachment points $c_{12}$ and $c_{21}$ to their nearest regular vertices in the direction of $c_{2}$ and $c_{1}$ respectively, resulting in a diangle of exceedance $e+1$, and the weight $1/t^{2}$ is needed to unweigh these regular vertices. ∎
Proof of Theorem A.1. .
We apply Proposition A.2, discussing separately the different cases that appear there. Let us first consider the case $a+b+c\in\mathbb{Z}_{>0}$: we may write
$T^{(0)}_{2a,2b,2c}$ $=\frac{1}{t^{6}}(R^{e(D_{1})}X)(R^{e(D_{2})}X)(R^{e(D_{3})}X)Y^{2}+\frac{R^{e(D_{1})+1}X}{t^{2}}\frac{R^{e(D_{2})+1}X}{t^{2}}\frac{R^{e(D_{3})+1}X}{t^{2}}\widetilde{Y}^{2}$
$=R^{a+b+c}\frac{X^{3}(Y^{2}+R^{3}\widetilde{Y}^{2})}{t^{6}}\,.$ (112)
Here, the term in $Y^{2}$ (resp. $\widetilde{Y}^{2}$) corresponds to the the situation where both $T_{1}$ and $T_{2}$ are even (resp. odd) triangles, and the division by $t^{6}$ in the first term arises from the identification between the vertices incident to the attachment points. Note that the second equality holds both for type I and type II, thanks to (111).
The formula (112) remains valid in the case $a=b=c=0$, except that we have to remove a spurious term $t^{-1}$ corresponding to the contribution of the quintuple whose all elements are reduced to the vertex map.
Let us now consider the case $a+b+c\in\mathbb{Z}_{\geq 0}+\frac{1}{2}$: then the discussion differs slightly between type I and type II. For type I, we may write
$T^{(0)}_{2a,2b,2c}=\frac{1}{t^{3}}\frac{R^{e(D_{1})+\frac{1}{2}}X}{t}\frac{R^{e(D_{2})+\frac{1}{2}}X}{t}\frac{R^{e(D_{3})+\frac{1}{2}}X}{t}(2Y\widetilde{Y})=R^{a+b+c-\frac{1}{2}}\frac{2R^{2}X^{3}Y\widetilde{Y}}{t^{6}}\,.$ (113)
Here, the factor $2Y\widetilde{Y}$ enumerates the pairs $(T_{1},T_{2})$ of different parities. Note that in this situation, all diangles have exactly one regular attachment point, and the factor $1/t^{3}$ arises from the identification between the vertices at these attachment points. For type II, assuming without loss of generality that $b\geq a+c$ as in Proposition A.2, we may write
$T^{(0)}_{2a,2b,2c}$ $=\frac{1}{t^{3}}\frac{R^{e(D_{1})+\frac{1}{2}}X}{t}\left((R^{e(D_{2})}X)\frac{R^{e(D_{3})+1}X}{t^{2}}Y\widetilde{Y}+\frac{R^{e(D_{2})+1}X}{t^{2}}(R^{e(D_{3})}X)\widetilde{Y}Y\right)$
$=R^{a+b+c-\frac{1}{2}}\frac{2R^{2}X^{3}Y\widetilde{Y}}{t^{6}}\,.$ (114)
Here, the first (resp. second) term corresponds to the case where $T_{1}$ is even (resp. odd), so that $D_{2}$ has two (resp. zero) regular attachment points, $D_{3}$ has zero (resp. two) regular attachment points. In all cases, $D_{1}$ has exactly one regular attachment point.
Gathering all these cases together, we arrive at the unified formula
$T^{(0)}_{2a,2b,2c}+t^{-1}\delta_{a+b+c=0}=\begin{cases}R^{a+b+c-1}\frac{RX^{3}(Y^{2}+R^{3}\widetilde{Y}^{2})}{t^{6}}&\text{if $a+b+c\in\mathbb{Z}_{\geq 0}$,}\\ R^{a+b+c-\frac{1}{2}}\frac{2R^{2}X^{3}Y\widetilde{Y}}{t^{6}}&\text{if $a+b+c\in\mathbb{Z}_{\geq 0}+\frac{1}{2}$.}\end{cases}$ (115)
We conclude by noting that
$\frac{\partial R}{\partial t}=T^{(0)}_{1,1,0}=\frac{RX^{3}(Y^{2}+R^{3}\widetilde{Y}^{2})}{t^{6}}\quad\mbox{and}\quad\frac{\partial S}{\partial t}=T^{(0)}_{1,0,0}=\frac{2R^{2}X^{3}Y\widetilde{Y}}{t^{6}}\,,$ (116)
since $R$ (resp. $S$) counts planar maps with two distinguished boundaries both of length $1$ (resp. one of length $1$ and the other of length 00): differentiating with respect to $t$ has the effect of adding a third boundary of length 00, and boundaries of lenghts 00 or $1$ are tautologically tight. ∎
We now state a variant of Theorem A.1 allowing to count tight pairs of pants with some boundaries strictly tight, under certain assumptions.
Proposition A.4.
Let $a$, $b$ and $c$ be nonnegative integers or half-integers, and let $T^{(0)}_{2a,2b|2c}$ (resp. $T^{(0)}_{2a|2b,2c}$) denote the generating function of planar maps with three labeled distinct tight boundaries of lengths $2a,2b,2c$, the third (resp. the second and the third) being strictly tight, where we attach a weight $t$ per vertex different from a boundary-vertex and not incident to the third boundary (resp. to the second nor to the third boundary) and, for all $k\geq 1$, a weight $t_{k}$ per inner face of degree $k$.
Then, for $a+b>c$, we have
$T^{(0)}_{2a,2b|2c}=\begin{cases}R^{a+b-c-1}\frac{\partial R}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}$,}\\ R^{a+b-c-\frac{1}{2}}\frac{\partial S}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}-\frac{1}{2}$,}\end{cases}$ (117)
while, for $a>b+c$, we have
$T^{(0)}_{2a|2b,2c}=\begin{cases}R^{a-b-c-1}\frac{\partial R}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}$,}\\ R^{a-b-c-\frac{1}{2}}\frac{\partial S}{\partial t}&\text{if $a+b+c\in\mathbb{Z}_{>0}-\frac{1}{2}$.}\end{cases}$ (118)
Observe that, in these two expressions, the assumptions $a+b>c$ and $a>b+c$ are respectively crucial to avoid having terms involving negative powers of $t$. It would be interesting to obtain formulas for $T^{(0)}_{2a,2b|2c}$ and $T^{(0)}_{2a|2b,2c}$ without these assumptions, but this is beyond the scope of this paper. Note that, whenever $a>b+c$, the boundaries of lengths $b,c$ cannot touch each other.
Proof of Proposition A.4.
In view of Theorem A.1, it suffices to establish that
$T^{(0)}_{2a,2b,2c}=\begin{cases}R^{2c}T^{(0)}_{2a,2b|2c}&\text{if $a+b>c$,}\\ R^{2b+2c}T^{(0)}_{2a|2b,2c}&\text{if $a>b+c$.}\\ \end{cases}$ (119)
Let us first assume $a+b>c$ and establish the formula for $T^{(0)}_{2a,2b|2c}$. If $c=0$ then there is nothing to prove, as a boundary-vertex is always considered strictly tight. Assuming $c>0$, let $M$ be a planar map with three labeled distinct tight boundaries of lengths $2a,2b,2c$. The third boundary is a face, which we denote by $F_{3}$. We then cut $M$ along the innermost minimal closed path $C$ homotopic to $F_{3}$. This path can be precisely defined as the maximal element of the lattice $\mathcal{C}_{\mathrm{min}}^{(3)}$, discussed in [triskell, Section 6.1]. We will justify in Lemma A.5 below that $C$ is simple, hence by cutting we obtain two pieces: a planar map with three tight boundaries of lengths $2a,2b,2c$, the third being strictly tight, and a planar map with two tight boundaries both of length $2c$. Moreover, the boundary resulting from this cutting operation can be canonically rooted: indeed, since planar maps with three boundaries have a trivial automorphism group, we may canonically select a distinguished corner incident to $C$. This decomposition is clearly bijective. Turning to generating functions, we decide to assign the weight $t$ for vertices on $C$ to the second map. It is known [irredmaps, Section 9.3] that planar maps with two tight boundaries both of length $2c$, one of which is rooted, have generating function $R^{2c}$. We conclude that $T^{(0)}_{2a,2b,2c}=R^{2c}T^{(0)}_{2a,2b|2c}$ as wanted.
For $a>b+c$, we may perform a similar decomposition with the boundary of length $2b$, giving the relation $T^{(0)}_{2a,2b|2c}=R^{2b}T^{(0)}_{2a|2b,2c}$. ∎
Let us now justify that the closed path $C$ used in the above proof is simple:
Lemma A.5.
Let $M$ be a planar map with three labeled distinct tight boundaries of lengths $\ell_{1},\ell_{2},\ell_{3}$. If $\ell_{1}+\ell_{2}>\ell_{3}>0$ then every closed path on $M$ of length $\ell_{3}$ which is homotopic to the third boundary is simple.
Proof.
Let $C$ be a closed path on $M$ of length $\ell>0$. We may decompose $C$ into a (finite) sequence of simple closed paths $\hat{C}=(C_{1},C_{2},\ldots,C_{n})$ with positive lengths, for instance by doing a “loop-erasure” (in this construction, a closed path reduced to a single edge followed once in both ways is regarded as simple).
To be explicit, choose an arbitrary vertex $v_{0}$ on $C$ and an orientation, and let us denote by $v_{0},e_{1},v_{1},e_{2},v_{2},\ldots,e_{\ell},v_{\ell}=v_{0}$ the successive vertices and edges of $M$ visited by $C$. We then apply the following iterative procedure. Initialize $\gamma$ as the path reduced to $v_{0}$, and $\hat{C}$ as an empty sequence. Then, for every step $i$ from $1$ to $\ell$, append the edge $e_{i}$ and $v_{i}$ to $\gamma$:
  • if $\gamma$ is simple, proceed to the next step,
  • if $\gamma$ is not simple, then it necessarily consists of a simple path $\gamma^{\prime}$ from $v_{0}$ to $v_{i}$ (possibly of length zero if $v_{i}=v_{0}$) and of a simple closed path $C^{\prime}$ with positive length: set $\gamma=\gamma^{\prime}$, append $C^{\prime}$ to the sequence $\hat{C}$, and proceed to the next step.
At the last step, the path $\gamma$ is reduced to $v_{0}=v_{\ell}$, and $\hat{C}$ is the sequence we are looking for. There is an interesting combinatorial structure of heap underlying this decomposition, see e.g. [Viennot1986, Proposition 6.3], but for our purposes we need only remark that the sum of the lengths of the $C_{i}$’s is $\ell$.
Now, let us assume that $C$ is homotopic to the third boundary. We claim that at least one of the following assertions is true:
  • (i)
    there exists $i$ such that $C_{i}$ is homotopic to the third boundary,
  • (ii)
    there exists $i,j$ such that $C_{i}$ and $C_{j}$ are homotopic to the first and second boundary, respectively.
Indeed, each $C_{i}$ is either contractible or homotopic to one of the boundaries, as we justify in Lemma A.7 below. It is not possible that all $C_{i}$ are contractible or homotopic to the first boundary, as otherwise $C$ would be homotopic to a multiple of the latter. Similarly for the second boundary. Hence, either (i) or (ii) must hold.
Let us now assume $\ell=\ell_{3}$. As the boundaries are assumed tight, in case (i) we find that $\hat{C}$ consists of a single element ($n=1$) so that $C=C_{1}$ is simple. In case (ii) $C$ would have length at least $\ell_{1}+\ell_{2}$, but this is excluded by the assumption $\ell_{1}+\ell_{2}>\ell_{3}$. ∎
Remark A.6.
In the case where $\ell_{1}$, say, vanishes, Lemma A.5 and the above proof still hold, upon using the convention discussed in Section 1.2 and [triskell, Figure 2] of replacing the boundary-vertex by a small circle made of edges of length zero. Observe that the boundary-vertex cannot be incident to the boundary-face of length $\ell_{3}$, as otherwise following the contour of that face but circumventing the boundary-vertex in the other direction would yield a path of length $\ell_{3}$ that is homotopic to the boundary-face of length $\ell_{2}$, contradicting the tightness assumption.
The following fact, used in the proof of Lemma A.5, seems well-known in some communities, but we rederive it for convenience.
Lemma A.7.
Let $M$ be a planar map with three boundaries. Then, every simple closed path on $M$ is either contractible or homotopic to one of the boundaries.
Proof.
We may draw $M$ in the plane, with the puncture of one of the boundaries sent to infinity. Let $C$ be a simple closed path on $M$. By the Jordan-Schoenflies theorem, $C$ delimits a bounded domain $D$ homeomorphic to a disk, which may contain zero, one or two punctures. If $D$ contains zero puncture, then $C$ is contractible. If $D$ contains one puncture, then $C$ is homotopic to the corresponding boundary. Finally, if $D$ contains two punctures, then $C$ is homotopic to the boundary corresponding to the puncture at infinity. ∎
Appendix B Quasi-polynomiality in the general case: induction step
This appendix and the next one are devoted to the proof of Proposition 4.8 by induction on $n$. Here, we check that if the statement of the proposition is true for $n$ boundaries, then it is also true for $n+1$ boundaries. We proceed in three steps.
B.1 A discrete integration lemma for polynomials
We first record an elementary discrete integration lemma. It is certainly well-known, but the derivation is a nice variation on the Bernoulli-Faulhaber formulas so we give a full proof.
Lemma B.1.
Fix an integer $k\geq 0$. Then the sums
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}m^{2k+1}+\frac{\ell^{2k+1}}{2}\,,\qquad\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}\,,$ (120)
are even polynomials in the even integer variable $\ell$, and the sums
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}m^{2k+1}\,,\qquad\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}+\frac{\ell^{2k+1}}{2}\,,$ (121)
are even polynomials in the odd integer variable $\ell$.
Proof.
First suppose that $\ell=2n$ is even. By the Bernoulli-Faulhaber formula
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}m^{2k+1}+\frac{\ell^{2k+1}}{2}$ $=2^{2k+1}\left(\sum_{q=1}^{n}q^{2k+1}-n^{2k+1}+\frac{n^{2k+1}}{2}\right)$ (122)
$=2^{2k+1}\left(\frac{1}{2k+2}\sum_{i=0}^{2k+1}\binom{2k+2}{i}B_{i}n^{2k+2-i}-\frac{n^{2k+1}}{2}\right)\,.$ (123)
Since $B_{1}=1/2$ and all other Bernoulli numbers with odd indices vanish, we obtain that this is a polynomial in $n^{2}=(\ell/2)^{2}$. Similarly, a simple consequence of the Bernoulli-Faulhaber formula, subtracting the terms with even indices, is that
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}=\frac{2^{2k+2}}{k+1}\sum_{i=0}^{2k+1}\binom{2k+2}{i}B_{i}n^{2k+2-i}(2^{1-i}-1)$ (124)
and we see that the term involving $B_{1}$ vanishes. Using again the vanishing properties of Bernoulli numbers of odd indices, this is a polynomial in $n^{2}=(\ell/2)^{2}$.
We now study the case where $\ell=2n-1$ is odd. Note that
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}m^{2k+1}=2^{2k+1}\sum_{0<q<n}q^{2k+1}$ (125)
which, by another formula of Faulhaber, is a polynomial in the variable $n(n-1)=(\ell^{2}-1)/4$, hence an even polynomial in $\ell$. Finally, simply note that, since $\ell$ is odd,
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}+\frac{\ell^{2k+1}}{2}=\sum_{0<m\leq\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}-\frac{\ell^{2k+1}}{2}=\frac{1}{2}\left(\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}+\sum_{0<m\leq\ell\atop m\in 2\mathbb{Z}+1}m^{2k+1}\right)\,,$ (126)
since the last term is the average of the two first equal terms. By a final use of the Bernoulli-Faulhaber formula, this quantity if of the form $(P(n)+P(n+1))$ where $P$ is a polynomial and $22n-1=\ell-$ 2, so it can be put in the form $P((\ell-1)/2)+P((\ell+1)/2)$, which is an even polynomial in $\ell$ by Newton’s formula. ∎
Corollary B.2.
Let $P\in\mathbb{Q}[x]$ be a polynomial. Then
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mP(m^{2})+\frac{\ell}{2}P(\ell^{2})\,,\qquad\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mP(m^{2})\,,$ (127)
are even polynomials of the even integer variable $\ell$, and
$\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mP(m^{2})\,,\qquad\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mP(m^{2})+\frac{\ell}{2}P(\ell^{2})\,,$ (128)
are even polynomials of the odd integer variable $\ell$.
B.2 Quasi-polynomiality for tight boundary insertions
Next, we prove the following result on the boundary insertion operator $D_{\ell}$.
Proposition B.3.
For every $i\geq 0$, the quantities
$R^{-\frac{\ell}{2}}\frac{D_{\ell}R^{(i)}}{R}\,\quad,\qquad R^{-\frac{\ell}{2}}\frac{D_{\ell}S^{(i)}}{\sqrt{R}}$ (129)
are parity-dependent quasi-polynomials in the variable $\ell^{2}$ whose coefficients are polynomial functions of $R^{(j)}/R,S^{(j)}/\sqrt{R},1\leq j\leq i+1$.
Proof.
We prove this statement by induction on $i$, using the fact that for every $\ell\geq 0$,
$D_{\ell}R^{(i)}=T^{(0)}_{\ell,1,1,0_{i}}\,,\qquad D_{\ell}S^{(i)}=T^{(0)}_{\ell,1,0_{i+1}}\,,$ (130)
where we use the simplifying notation $00_{i}=0,\ldots,$ 0 for a string of $i$ consecutive zeros. We initialize the induction at $i=0$ by noting that, by our formula (109) for tight pairs of pants,
$D_{\ell}R=R^{\frac{\ell}{2}+1}\times\begin{cases}\displaystyle{\frac{R^{\prime}}{R}}&\text{if $\ell$ is even}\\ \displaystyle{\frac{S^{\prime}}{\sqrt{R}}}&\text{ if $\ell$ is odd,}\end{cases}\qquad D_{\ell}S=R^{\frac{\ell+1}{2}}\times\begin{cases}\displaystyle{\frac{S^{\prime}}{\sqrt{R}}}&\text{if $\ell$ is even}\\ \displaystyle{\frac{R^{\prime}}{R}}&\text{ if $\ell$ is odd.}\end{cases}$ (131)
Suppose that the statement holds up to some given value of $i$: namely, we may find polynomials $P_{j,2},P_{j,3},Q_{j,1},Q_{j,2}$ in $\mathbb{Q}[x_{1},\ldots,x_{j+1},y_{1},\ldots,y_{j+1},u]$ (the second index will refer to the number of boundary-faces of odd degree in the expressions below) such that, for $0\leq j\leq i$, and letting $\epsilon$ be 00 if $\ell$ is even and $1$ if $\ell$ is odd, we have
$\begin{split}D_{\ell}R^{(j)}&=R^{\frac{\ell}{2}+1}P_{j,2+\epsilon}\left(\frac{R^{(\leq j+1)}}{R},\frac{S^{(\leq j+1)}}{\sqrt{R}},\ell^{2}\right)\,,\\ D_{\ell}S^{(j)}&=R^{\frac{\ell+1}{2}}Q_{j,1+\epsilon}\left(\frac{R^{(\leq j+1)}}{R},\frac{S^{(\leq j+1)}}{\sqrt{R}},\ell^{2}\right)\,.\end{split}$ (132)
Here, we use the shorthand notation $R^{(\leq j+1)},S^{(\leq j+1)}$ for the variables $(R^{(k)},1\leq k\leq j+1)$ and $(S^{(k)},1\leq k\leq j+1)$.
Now, to prove the statement at rank $i+1$, we write
$D_{\ell}R^{(i+1)}$ $=T^{(0)}_{\ell,1,1,0_{i+1}}=\frac{\partial}{\partial t}T^{(0)}_{\ell,1,1,0_{i}}+\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,1,0_{i}}\,,$ (133)
$D_{\ell}S^{(i+1)}$ $=T^{(0)}_{\ell,1,0_{i+2}}=\frac{\partial}{\partial t}T^{(0)}_{\ell,1,0_{i+1}}+\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,0_{i+1}}\,,$ (134)
where the second equality on each line is a consequence of Proposition 3.3. Let us first analyse the right-hand side of (133). Its first term is equal to
$\frac{\partial}{\partial t}T^{(0)}_{\ell,1,1,0_{i}}$ $=\left(\frac{\ell}{2}+1\right)R^{\frac{\ell}{2}+1}\frac{R^{\prime}}{R}P_{i,2+\epsilon}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)$ (135)
$+R^{\frac{\ell}{2}+1}\sum_{k=1}^{i+1}\left(\frac{R^{(k)}}{R}\right)^{\prime}\frac{\partial P_{i,2+\epsilon}}{\partial x_{k}}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)$ (136)
$+R^{\frac{\ell}{2}+1}\sum_{k=1}^{i+1}\left(\frac{S^{(k)}}{\sqrt{R}}\right)^{\prime}\frac{\partial P_{i,2+\epsilon}}{\partial y_{k}}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)\,.$ (137)
Note that $(R^{(k)}/R)^{\prime}=(R^{(k+1)}/R)-(R^{(k)}/R)(R^{\prime}/R)$ and $(S^{(k)}/\sqrt{R})^{\prime}=(S^{(k+1)}/\sqrt{R})-(1/2)(S^{(k)}/\sqrt{R})(R^{\prime}/R)$, and therefore, the last two sums are quasi-polynomial in $\ell^{2}$ with coefficients in $\mathbb{Q}[R^{(\leq i+2)}/R,S^{(\leq i+2)}/\sqrt{R}]$. In order to study the last sum in (133), it is easier to split cases.
Assume that $\ell$ is even ($\epsilon=0$).
The last sum in (133) then equals
$\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,1,0_{i}}$ $=\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mR^{\frac{\ell-m}{2}}\frac{R^{\prime}}{R}R^{\frac{m}{2}+1}P_{i,2}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)$ (138)
$+\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mR^{\frac{\ell-m}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{m}{2}+1}P_{i,3}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)\,.$ (139)
By Corollary B.2, the term (139), and the sum of the term (138) and (135) (for $\epsilon=0$) is of the wanted form of $R^{\frac{\ell}{2}+1}$ times an even polynomial in the variable $\ell^{2}$, with coefficients in $\mathbb{Q}[R^{(\leq i+1)}/R,S^{(\leq i+1)}/\sqrt{R}]$. Putting things together, we see that (133) indeed yields an even polynomial expression in the even variable $\ell$.
Assume that $\ell$ is odd ($\epsilon=1$).
This time, we have
$\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,1,0_{i}}$ $=\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mR^{\frac{\ell-m}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{m}{2}+1}P_{i,2}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)$ (140)
$+\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mR^{\frac{\ell-m}{2}}\frac{R^{\prime}}{R}R^{\frac{m}{2}+1}P_{i,3}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)\,.$ (141)
We then apply Corollary B.2 to the term (140) on the one hand, and to the sum of the term (141) and (135) (for $\epsilon=1$) on the other hand. These are of the wanted form of $R^{\frac{\ell}{2}+1}$ times an even polynomial in $\ell^{2}$, with coefficients in $\mathbb{Q}[R^{(\leq i+1)}/R,S^{(\leq i+1)}/\sqrt{R}]$. Putting things together, we see that (133) indeed yields an even polynomial expression in the odd variable $\ell$. This completes the induction step for $D_{\ell}R^{(i+1)}$.
We now analyse of the right-hand side of (134) in a similar way. Its first term is equal to
$\frac{\partial}{\partial t}T^{(0)}_{\ell,1,0_{i+1}}$ $=\frac{\ell+1}{2}\cdot R^{\frac{\ell+1}{2}}\frac{R^{\prime}}{R}Q_{i,1+\epsilon}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)$ (142)
$+R^{\frac{\ell+1}{2}}\sum_{k=1}^{i+1}\left(\frac{R^{(k)}}{R}\right)^{\prime}\frac{\partial Q_{i,1+\epsilon}}{\partial x_{k}}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)$ (143)
$+R^{\frac{\ell+1}{2}}\sum_{k=1}^{i+1}\left(\frac{S^{(k)}}{\sqrt{R}}\right)^{\prime}\frac{\partial Q_{i,1+\epsilon}}{\partial y_{k}}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},\ell^{2}\right)$ (144)
Again, the last two sums are quasi-polynomial in the variable $\ell^{2}$ with coefficients in $\mathbb{Q}[R^{(\leq i+2)}/R,S^{(\leq i+2)}/\sqrt{R}]$. To analyse the last sum in (134) we again split cases. Assume that $\ell$ is even ($\epsilon=0$). Then, we have
$\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,0_{i+1}}$ $=\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mR^{\frac{\ell-m}{2}}\frac{R^{\prime}}{R}R^{\frac{m+1}{2}}Q_{i,1}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)$ (145)
$+\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mR^{\frac{\ell-m}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{m+1}{2}}Q_{i,2}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)\,.$ (146)
We apply Corollary B.2 to the term (146), and to the sum of the term (145) and (142) (for $\epsilon=0$), showing that they are of the wanted form of $R^{\frac{\ell+1}{2}}$ times an even polynomial in the variable $\ell^{2}$, with coefficients in $\mathbb{Q}[R^{(\leq i+1)}/R,S^{(\leq i+1)}/\sqrt{R}]$. Finally, assume that $\ell$ is odd ($\epsilon=1$). This time, we have
$\sum_{0<m<\ell}mT^{(0)}_{\ell,0|m}T^{(0)}_{m,1,0_{i+1}}$ $=\sum_{0<m<\ell\atop m\in 2\mathbb{Z}}mR^{\frac{\ell-m}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{m+1}{2}}Q_{i,1}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)$ (147)
$+\sum_{0<m<\ell\atop m\in 2\mathbb{Z}+1}mR^{\frac{\ell-m}{2}}\frac{R^{\prime}}{R}R^{\frac{m+1}{2}}Q_{i,2}\left(\frac{R^{(\leq i+1)}}{R},\frac{S^{(\leq i+1)}}{\sqrt{R}},m^{2}\right)\,.$ (148)
We then apply Corollary B.2 to the term (147) on the one hand, and to the sum of the term (148) and (142) (for $\epsilon=1$) on the other hand. These are of the wanted form of $R^{\frac{\ell}{2}+1}$ times an even polynomial in the variable $\ell^{2}$, with coefficients in $\mathbb{Q}[R^{(\leq i+1)}/R,S^{(\leq i+1)}/\sqrt{R}]$. Putting things together, we see that (134) indeed yields an even polynomial expression in the odd variable $\ell$. This completes the induction step for $D_{\ell}S^{(i+1)}$, and the proof of Proposition B.3. ∎
From the fact that $D_{m}$ is a differential operator, note that
$D_{\ell}\left(\frac{R^{(i)}}{R}\right)=\frac{D_{\ell}R^{(i)}}{R}-\frac{R^{(i)}}{R}\frac{D_{\ell}R}{R}\,,\qquad D_{\ell}\left(\frac{S^{(i)}}{\sqrt{R}}\right)=\frac{D_{\ell}S^{(i)}}{\sqrt{R}}-\frac{1}{2}\frac{S^{(i)}}{\sqrt{R}}\frac{D_{\ell}R}{R}$ (149)
are of the form of $R^{\ell/2}$ times a parity-dependent quasi-polynomial in $\ell^{2}$, and with coefficients in $\mathbb{Q}[R^{(\leq i+1)}/R,S^{(\leq i+1)}/\sqrt{R}]$.
B.3 End of the proof of the induction step
Let us now assume that the statement of Proposition 4.8 is true for $(g,n)$, we will show that it is then true for $(g,n+1)$. Let us introduce the shorthand notation
$d_{g,n}:=3g-2+n.$ (150)
By the induction hypothesis, for every subset $I$ of $\{1,\ldots,n\}$ there exists
$P^{(g)}_{n,I}\in\mathbb{Q}(x_{1},\ldots,x_{d_{g,n}},y_{1},\ldots,y_{d_{g,n}})[u_{1},\ldots,u_{n}]\,,$ (151)
such that, whenever the $\ell_{i}$ with $i\in I$ are odd integers, and the $\ell_{i}$ with $i\notin I$ are even integers, we have
$\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}=P^{(g)}_{n,I}\left(\frac{R^{(\leq d_{g,n})}}{R},\frac{S^{(\leq d_{g,n})}}{\sqrt{R}},\ell_{1}^{2},\ldots,\ell_{n}^{2}\right)-c(g,n)t^{2-2g-n}\delta_{\ell_{1}+\cdots+\ell_{n},0}\,.$ (152)
The $P^{(g)}_{n,I}$ correspond to the family of polynomials associated with the parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$, but the extra variables $x_{1},\ldots,x_{d_{g,n}},y_{1},\ldots,y_{d_{g,n}}$ are useful to keep track of the dependency in the derivatives of $R$ and $S$. In what follows, to lighten notation, we omit the mention of these extra variables, except when we differentiate with respect to them.
Recall that Propositions 3.3 and 3.5 assert that
$T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}=D_{\ell_{n+1}}T^{(g)}_{\ell_{1},\ldots,\ell_{n}}+\sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}m_{i}\,T^{(0)}_{\ell_{i},\ell_{n+1}|m_{i}}T^{(g)}_{\ell_{1},\ldots,m_{i},\ldots,\ell_{n}}$ (153)
with $D_{0}=\frac{\partial}{\partial t}$ and $D_{m}$ as in (61) for $m>0$. We plug (152) into the above relation. Since $D_{\ell_{n+1}}$ is a differential operator that annihilates the variable $t$ except when $\ell_{n+1}=0$, the first term is equal to
$D_{\ell_{n+1}}\left(R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\right)=\frac{\ell_{1}+\cdots+\ell_{n}}{2}\cdot R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\frac{D_{\ell_{n+1}}R}{R}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\\ +R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\sum_{r=1}^{d_{g,n}}\left(D_{\ell_{n+1}}\left(\frac{R^{(r)}}{R}\right)\frac{\partial P^{(g)}_{n,I}}{\partial x_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})+D_{\ell_{n+1}}\left(\frac{S^{(r)}}{\sqrt{R}}\right)\frac{\partial P^{(g)}_{n,I}}{\partial y_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\right)\\ -c(g,n+1)t^{2-2g-(n+1)}\delta_{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1},0}$ (154)
where we set $c(g,n+1)=(2-2g-n)c(g,n)$. By Proposition B.3 and its consequence discussed around (149), the second term of (154) is of the wanted form $R^{\frac{\ell_{1}+\cdots+\ell_{n+1}}{2}}$ times a quasi-polynomial in $\ell_{1}^{2},\ldots,\ell_{n+1}^{2}$. We then split the second term of (153) into pieces.
Assume first that that $\ell_{n+1}$ is even, in which case, the quantity $D_{\ell_{n+1}}R/R$ appearing in the first term of (154) equals $R^{\ell_{n+1}/2}R^{\prime}/R$. The last term of (153) is then
$\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{R^{\prime}}{R}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (155)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}+1}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I\cup\{i\}}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (156)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}+1}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I\setminus\{i\}}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (157)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}+1}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}+1}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{R^{\prime}}{R}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})\,.$ (158)
By the discrete integration Corollary B.2, we see that (156) and (157) are polynomial expressions in their respectively even and odd variables $\ell_{i}^{2}$, while we should combine (155) and (158) with the corresponding first terms of (154) (which is made possible by the fact that $D_{\ell_{n+1}}$ has a factor $R^{\prime}/R$) to get similar polynomial expressions.
If, on the other hand, $\ell_{n+1}$ is odd, then $D_{\ell_{n+1}}R/R$ equals $R^{\ell_{n+1}/2}S^{\prime}/\sqrt{R}$. The last term of (153) is then
$\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (159)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}+1}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{R^{\prime}}{R}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I\cup\{i\}}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (160)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}+1}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{R^{\prime}}{R}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I\setminus\{i\}}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})$ (161)
$+\sum_{i=1\atop\ell_{i}\in 2\mathbb{Z}+1}^{n}\sum_{0<m_{i}<\ell_{i}\atop m_{i}\in 2\mathbb{Z}+1}m_{i}R^{\frac{\ell_{i}+\ell_{n+1}-m_{i}}{2}}\frac{S^{\prime}}{\sqrt{R}}R^{\frac{\ell_{1}+\cdots+m_{i}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,m_{i}^{2},\ldots,\ell_{n}^{2})\,.$ (162)
By the discrete integration Corollary B.2, we see that (160) and (161) are polynomial expressions in their respectively even and odd variables $\ell_{i}^{2}$, while we should combine (159) and (162) with the corresponding first terms of (154) to get similar polynomial expressions.
From this discussion we conclude that the statement of Proposition 4.8 holds for $(g,n+1)$, which concludes the proof of the induction step.
Appendix C Quasi-polynomiality in the general case: initialisation
We now initialise the induction for proving Proposition 4.8: for $g=0$, $g=1$ and $g\geq 2$ we should respectively show that the statement holds for $n=3$, $n=1$ and $n=0$.
That the statement holds in the planar case $(g,n)=(0,3)$ is a direct consequence of Theorem A.1 and of the relation (23) between the $T$’s and the $\tau$’s. To treat the case of higher genera, we will need some results from the theory of topological recursion presented in [Eynard2016], and more precisely, we will need the expression of the generating function of maps in genus $g\geq 1$ based on the so-called method of moments [Ambjoern1995]. Our approach can be seen as an extension of the analysis of Budd [budd2020irreducible, Section 3.4] to the non bipartite case (but only considering the non irreducible case). Recall that $F^{(g)}=F^{(g)}_{\varnothing}=T^{(g)}_{\varnothing}=\tau^{(g)}_{\varnothing}$ denote the generating series of maps of genus $g$ without boundaries.
C.1 Genus one
Let us first deal with the case $g=1$. Our main input will be the following.
Theorem C.1.
The generating function for maps of genus one without boundary is given by
$F^{(1)}=\frac{1}{24}\ln\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)+\frac{\ln(t)}{12}\,.$ (163)
From this, we can easily deduce Proposition 4.8 in the case $(g,n)=(1,1)$. Indeed, note that by Propositions 3.3 and 3.5, we have $T^{(1)}_{\ell}=D_{\ell}F^{(1)}$, which equals
$T^{(1)}_{\ell}=\frac{\frac{R^{\prime}}{R}D_{\ell}\left(\frac{R^{\prime}}{R}\right)-\frac{S^{\prime}}{\sqrt{R}}D_{\ell}\left(\frac{S^{\prime}}{\sqrt{R}}\right)}{12\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)}+\frac{\delta_{\ell,0}}{12t}$ (164)
and by Proposition B.3—see also (149)—this is of the wanted form of $R^{\ell/2}$ times a quasipolynomial in $\ell$ with quasiperiod $2$, whose coefficients are rational functions of $R^{(i)}/R,S^{(i)}/\sqrt{R},i\in\{1,2\}$, minus a pathological term for $\ell=0$.
Proof of Theorem C.1.
Let us recall Eynard’s notation from [Eynard2016] (Eynard puts a bound on the degrees, and excludes faces of degrees $1$ and $2$, we are going to ignore this). We set (p.58 and Theorem 3.1.2 p.62) $x(z)=S+\sqrt{R}(z+1/z)$ ($S=\alpha,R=\gamma^{2}$), and (pp.59-60, Definition 3.1.1 p.63 and Theorem 3.1.2 p.62)
$y(z)$ $=W_{1}^{(0)}(x(z))-\frac{V^{\prime}(x(z))}{2}$ (165)
$=-\frac{1}{2}\sum_{k\geq 1}u_{k}(z^{k}-z^{-k})\,,$ (166)
where
$u_{k}=S\delta_{k,0}+\sqrt{R}\delta_{k,1}-\sum_{i\geq k+1}t_{i}\sum_{j=k}^{[(i+k)/2]-1}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\,.$ (167)
Note that the latter is related to the generating function $V_{k-1}$ for $(k-1)$-slices introduced in [irredmaps]: for $k\geq 2$, we have $V_{k-1}=-R^{k/2}u_{k}$. On the other hand, we have $u_{0}=0$ and $u_{1}=t/\sqrt{R}$, which amounts to (3). We define the 00-th order moments by
$M_{+,0}=\frac{-y^{\prime}(1)}{\sqrt{R}}\,,\qquad M_{+,0}=\frac{-y^{\prime}(-1)}{\sqrt{R}}.$ (168)
Now, [Eynard2016, Theorem 3.4.6] expresses the generating function of toric maps in terms of $M_{\pm,0}$ as
$F^{(1)}=-\frac{1}{24}\ln\left(\frac{\gamma^{2}y^{\prime}(1)y^{\prime}(-1)}{t^{2}}\right)=-\frac{1}{24}\ln\left(\frac{R^{2}M_{+,0}M_{-,0}}{t^{2}}\right)\,,$ (169)
On the other hand, [Eynard2016, Theorem 3.3.4] gives
$\begin{split}\omega_{3}^{(0)}(z_{1},z_{2},z_{3})=&\frac{1}{2RM_{+,0}}\frac{1}{(z_{1}-1)^{2}(z_{2}-1)^{2}(z_{3}-1)^{2}}\\ &-\frac{1}{2RM_{-,0}}\frac{1}{(z_{1}+1)^{2}(z_{2}+1)^{2}(z_{3}+1)^{2}}\,,\end{split}$ (170)
which we now know to count pairs of pants. Namely, this is
$\sum_{\ell_{1},\ell_{2},\ell_{3}\geq 1}\left(\frac{1}{2RM_{+,0}}+\frac{(-1)^{\ell_{1}+\ell_{2}+\ell_{3}}}{2RM_{-,0}}\right)\frac{\ell_{1}\ell_{2}\ell_{3}}{z_{1}^{\ell_{1}+1}z_{2}^{\ell_{2}+1}z_{3}^{\ell_{3}+1}}\,.$ (171)
Comparing with our formula for tight pairs of pants (109), we obtain
$\tau^{(0)}_{\ell_{1},\ell_{2},\ell_{3}}=\left\{\begin{split}\frac{1}{2RM_{+,0}}+\frac{1}{2RM_{-,0}}&=\frac{R^{\prime}}{R}\mbox{ if }\ell_{1}+\ell_{2}+\ell_{3}\in 2\mathbb{Z}\\ \frac{1}{2RM_{+,0}}-\frac{1}{2RM_{-,0}}&=\frac{S^{\prime}}{\sqrt{R}}\mbox{ if }\ell_{1}+\ell_{2}+\ell_{3}\in 2\mathbb{Z}+1\end{split}\right.,$ (172)
where $\ell_{1},\ell_{2},\ell_{3}>0$. This shows that
$\frac{1}{RM_{+,0}}=-\frac{1}{\gamma y^{\prime}(1)}=\frac{R^{\prime}}{R}+\frac{S^{\prime}}{\sqrt{R}}\,,\qquad\frac{1}{RM_{-,0}}=-\frac{1}{\gamma y^{\prime}(-1)}=\frac{R^{\prime}}{R}-\frac{S^{\prime}}{\sqrt{R}}\,.$ (173)
Plugging this into (169) yields Theorem C.1. ∎
C.2 Higher moments, higher genera
The generating functions for maps in higher genera are given by higher order moments, which are given ([Eynard2016, p.64]), for $h\geq 1$, by
$M_{\pm,h}=\frac{-1}{R^{(h+1)/2}M_{\pm,0}}\sum_{k\geq h+1}(\pm 1)^{k+h+1}u_{k}\binom{k+h}{2h+1}\,.$ (174)
We define the renormalized version
$\bar{M}_{\pm,h}=R^{(h+1)/2}M_{\pm,h}M_{\pm,0}=-\sum_{k\geq h+1}(\pm 1)^{k+h+1}u_{k}\binom{k+h}{2h+1}\,.$ (175)
One should note that [Eynard2016, Corollary 3.5.1] introduces the quantities $R^{h/2}M_{\pm,h}$ that are called “dimensionless”, we will see that these are indeed natural quantities to consider, due to the following result [^8] There seems to be an inaccuracy in the statement of [Eynard2016, Corollary 3.5.1]. Indeed, the latter is stated without the boundary term involving $t^{2-2g}$ only, which is however present in other similar statements such as [ibid., Theorems 3.4.3 and 3.4.9]. In these statements, the coefficient $\frac{B_{2g}}{2g(2-2g)}$ is indeed equal to $\chi(\mathcal{M}_{g,0})$ as wanted. It seems that this boundary term got omitted at some stage.:
Theorem C.2.
[Eynard2016, Corollary 3.5.1] There exists $\mathcal{P}_{g}\in\mathbb{Q}[x_{+},x_{-},(y_{+,h},y_{-,h},1\leq h\leq 3g-3)]$ that is homogeneous of degree $22g-$ 2 in the first two variables, and such that $\mathcal{P}_{g}(x_{+},x_{-},(\mu^{h},\mu^{h},1\leq h\leq 3g-3))$ is of degree $33g-$ 3 in $\mu$, and $c(g,0)\in\mathbb{Q}$, such that
$F^{(g)}=\mathcal{P}_{g}\left(\frac{1}{RM_{+,0}},\frac{1}{RM_{-,0}},(R^{h/2}M_{+,h},R^{h/2}M_{-,h},1\leq h\leq 3g-3)\right)-c(g,0)t^{2-2g}\,.$ (176)
We would like to show that—as for $M_{\pm,0}$, recall (173)—the “dimensionless” moments $R^{h/2}M_{\pm,h}$ are rational functions of $R^{(k)}/R,S^{(k)}/\sqrt{R}$ for $11\leq k\leq h+$ 1. Our goal will be to get rid of the variables $t_{i}$ appearing implicitly in the $u_{k}$ in the previous expression, and express this purely in terms of $R,S$ and their derivatives. Our approach is inspired by that of [budd2020irreducible, Section 3.4].
First, for $h\geq 1$, we rewrite, using (174) and (167),
$\bar{M}_{+,h}=\sum_{i\geq 2}t_{i}\sum_{k}\binom{k+h}{2h+1}\sum_{j}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\\ =\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}$ (177)
where the second expression comes from substituting the summation over $j$ to the summation over $l=i-1+k-2j$. To alleviate notations, we do not specify the summation ranges for $k,j,l$ as they are naturally enforced by the vanishing of the binomial and multinomial coefficients and by the requirement that their arguments are nonnegative integers. Note that, in the summation over $i$, the terms $i=2,\ldots,h+1$ have a vanishing contribution.
C.2.1 A binomial identity
We now state the following variant of [budd2020irreducible, Lemma 7].
Lemma C.3.
For any nonnegative integer $h$, there exist two polynomials $Q_{h}^{[0]}$ and $Q_{h}^{[1]}$ of degree $h+1$ such that, denoting by $\epsilon=\epsilon(i,l)\in\{0,1\}$ the parity of $i+l$, we have
$\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}=\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q_{h}^{[\epsilon]}\left(\frac{i-l-\epsilon}{2}\right)$ (178)
where we sum over nonnegative $k$ such that $k+\epsilon$ is odd.
Proof.
We rewrite the sum as
$\binom{i-1}{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-1-l}{\frac{i-1+k-l}{2}}$ (179)
Assume that $i-l$ is even, that is, $\epsilon=0$, so that the summation is over odd $k=2k^{\prime}+1$. Then, writing $2j^{\prime}=i-l$, this rewrites
$\binom{i-1}{l}\sum_{k^{\prime}}\binom{2k^{\prime}+h+1}{2h+1}\binom{2j^{\prime}-1}{j^{\prime}+k^{\prime}}=\binom{i-1}{l}\binom{2j^{\prime}-1}{j^{\prime}}Q_{h}(j^{\prime})\,,$ (180)
where $Q_{h}=Q_{h}(0,\cdot)$ is the polynomial of degree $h+1$ that appears in [budd2020irreducible, Lemma 7]. Alternatively, we can derive this lemma by the following independent argument. First note that, in the notation of [polytightmaps], we have
$\binom{x+h}{2h+1}=\frac{h!^{2}}{(2h+1)!}xp_{h}(x)\,,$ (181)
so that, using the notation
$\mathsf{A}_{j,k}=\frac{2k}{2j}\binom{2j}{j+k}\,,\quad j,k\in\frac{1}{2}\mathbb{Z}$ (182)
from [polytightmaps, Equation (3.2)], we can rewrite the left hand side of (180) in the form
$\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}2(k^{\prime}+1/2)p_{h}(2(k^{\prime}+1/2))\binom{2(j^{\prime}-1/2)}{j^{\prime}-1/2+k^{\prime}+1/2}\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}p_{h}(2(k^{\prime}+1/2))\,.$ (183)
The polynomial $p_{h}(2m)$ is of degree $h$ in $m^{2}$, where $m\in\mathbb{Z}+1/2$, so it can be expressed in the basis of the polynomials $\tilde{p}_{r}(m),0\leq r\leq h$ defined in [polytightmaps, Equation (2.18)], in the form $p_{h}(2m)=\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\tilde{p}_{r}(m)$, which, based on [polytightmaps, Equation (3.9)], yields the expression
$(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-2}{j^{\prime}-1}\\ =\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\,j^{\prime}\binom{j^{\prime}-1}{r}$ (184)
which, since $j^{\prime}=(i-l)/2$, is indeed of the wanted form
$\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}Q_{h}^{[0]}\left(\frac{i-l}{2}\right)\,,$ (185)
for some polynomial $Q^{[0]}_{h}$ of degree $h+1$.
For $\epsilon=1$, we argue in a similar (and slightly simpler) way. The summation is now over even $k=2k^{\prime}$, and we write $12j^{\prime}=i-l-$ 1, so that (179) rewrites, from [polytightmaps, Equation (3.2)]:
$\binom{i-1}{l}\sum_{k^{\prime}}\binom{2k^{\prime}+h}{2h+1}\binom{2j^{\prime}}{j^{\prime}+k^{\prime}}=2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{h}(2k^{\prime})\,.$ (186)
This time, we express the polynomial $p_{h}(2m)$ in the basis of the polynomials $p_{r}(m),0\leq r\leq h$, in the form $p_{h}(2m)=\sum_{r=0}^{h}\alpha_{h,r}p_{r}(k)$, which yields, now using [polytightmaps, Equation (3.7)], the expression
$2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\sum_{r=0}^{h}\alpha_{h,r}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\alpha_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-1}{j^{\prime}}\\ =\binom{i-1}{\frac{i-l-1}{2},\frac{i-l-1}{2},l}\frac{h!^{2}}{(2h+1)!}j^{\prime}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\,,$ (187)
which is again of the wanted form. ∎
C.2.2 Expressing the moments in terms of derivatives of $\mathbf{Z}$
Lemma C.3 and (177) allow to rewrite the moment $M_{+,h}$ as
$\bar{M}_{+,h}=\sum_{\epsilon\in\{0,1\}}\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q^{[\epsilon]}_{h}\left(\frac{i-l-\epsilon}{2}\right)\ ,$ (188)
and
$(-1)^{h}\bar{M}_{-,h}=\sum_{\epsilon\in\{0,1\}}(-1)^{\epsilon}\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q^{[\epsilon]}_{h}\left(\frac{i-l-\epsilon}{2}\right)\ ,$ (189)
Now, let us rewrite the equations (3) defining $R$ and $S$ in the more symmetric form
$\mathbf{Z}(\mathbf{U}(\mathbf{t}))=\mathbf{t}$ (190)
where $\mathbf{t}=(t_{0},t_{1})$, $t_{0}=t$, $\mathbf{U}=(R,S)$ and $\mathbf{Z}=(Z_{0},Z_{1})$ is the pair of bivariate series defined by
$Z_{0}(r,s)=r-\sum_{i\geq 2}t_{i}\sum_{l}\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}r^{\frac{i-l}{2}}s^{l}$ (191)
and
$Z_{1}(r,s)=s-\sum_{i\geq 2}t_{i}\sum_{l}\binom{i-1}{\frac{i-l-1}{2},\frac{i-l-1}{2},l}r^{\frac{i-l-1}{2}}s^{l}\,.$ (192)
Then we obtain that
$\bar{M}_{+,h}=R^{-1/2}Q^{[0]}_{h}(r\partial_{r})Z_{0}(r,s)|_{r=R,s=S}+Q^{[1]}_{h}(r\partial_{r})Z_{1}(r,s)|_{r=R,s=S}\,,$ (193)
and
$(-1)^{h}\bar{M}_{h,-}=R^{-1/2}Q^{[0]}_{h}(r\partial_{r})Z_{0}(r,s)|_{r=R,s=S}-Q^{[1]}_{h}(r\partial_{r})Z_{1}(r,s)|_{r=R,s=S}\,.$ (194)
By Leibniz’ formula, this means that the (renormalized) moments are (possibly with some factors $R^{-1/2}$) linear combinations of the expressions
$R^{k}\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(R,S)\,,\qquad 1\leq k\leq h+1\,,\quad\epsilon\in\{0,1\}.$ (195)
We now discuss the structure of these expressions.
C.2.3 Structure of derivatives of $\mathbf{Z}$
To this end, we need to understand better the partial derivatives of $R,S$ with respect to $t_{0},t_{1}$. By applying a simple induction argument based on Proposition B.3 (and noting that the differential operators $D_{\ell}$ for $\ell\in\{0,1\}$ are given by $D_{\ell}=\partial/\partial t_{\ell}$), we obtain the following statement.
Proposition C.4.
One has
$\frac{\partial^{a+b}R}{\partial t_{0}^{a}\partial t_{1}^{b}}=(a+b)!R^{\frac{b}{2}+1}P_{0}^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)$ (196)
and
$\frac{\partial^{a+b}S}{\partial t_{0}^{a}\partial t_{1}^{b}}=(a+b)!R^{\frac{b+1}{2}}P_{1}^{(a,b)}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq a+b\right)$ (197)
where $P^{(a,b)}_{0},P^{(a,b)}_{1}$ are polynomials of valuation [^9] In Eynard’s book, this would be coined as “homogeneous, where $R^{(k)}$ and $S^{(k)}$ are considered of degree $k$”. $a+b$, where a derivative of order $k$ counts for $k$ in the valuation. The normalization $(a+b)!$ is here to make things more convenient afterwards.
Note that these derivatives are of the form $T^{(0)}_{1,\ldots,1,0,\ldots 0}$ in both cases, with $a$ terms equal to 00 and $b+2$ terms equal to $1$ in the first case, and with $a+1$ terms equal to 00 and $b+1$ terms equal to $1$ in the second case, so in fact we have $P_{0}^{(a+1,b-1)}=P_{1}^{(a,b)}$.
At this point, we use the formula (212) for derivatives of inverses discussed in Appendix D, applied to $\mathbf{f}=\mathbf{U}$ and $\mathbf{g}=\mathbf{Z}$ (for $n=2$ in the notation therein)
$\frac{\partial^{k}Z_{\epsilon}}{\partial r^{k}}(\mathbf{U}(\mathbf{t}))=\sum_{\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{t}}\mathbf{U})^{-1}_{i_{v},j_{v}}\prod_{v\in\tau,k_{v}(\tau)>0}\frac{1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}U_{j_{v}}}{\partial t_{i_{v1}}\cdots\partial t_{i_{vk_{v}(\tau)}}}(\mathbf{t})$ (198)
where $\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$ is the family of planted Pólya trees $\tau$ with $k$ leaves, and where the bottom and top half-edges of the edge $e_{v}$ below the vertex $v\in\tau$ are labeled by elements $i_{v},j_{v}\in\{0,1\}$, in such a way that $i_{\varnothing}=\epsilon$, and $i_{v}=0$ for every leaf $v\in\tau$.
Let us look more closely at the contribution of a given tree $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$ to the preceding sum. Note that by (131), we have
$d_{\mathbf{t}}\mathbf{U}=\left(\begin{array}[]{cc}\frac{\partial R}{\partial t_{0}}&\frac{\partial R}{\partial t_{1}}\\ \frac{\partial S}{\partial t_{0}}&\frac{\partial S}{\partial t_{1}}\end{array}\right)=\left(\begin{array}[]{cc}R^{\prime}&RS^{\prime}\\ S^{\prime}&R^{\prime}\end{array}\right)$ (199)
We use the comatrix formula to compute
$(d_{\mathbf{t}}\mathbf{U}^{-1})_{i,j}=R^{-1-\frac{i}{2}+\frac{j}{2}}\frac{\mathcal{D}(i,j)}{\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}}\,,$ (200)
where $\mathcal{D}(i,j)$ is $R^{\prime}/R$ if $i+j$ is even, and $-S^{\prime}/\sqrt{R}$ otherwise. Hence, the contribution of edges of $\tau$—the first product in (198)—is
$\prod_{v\in V(\tau)}R^{-1-\frac{i_{v}}{2}+\frac{j_{v}}{2}}\frac{\mathcal{D}(i_{v},j_{v})}{\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}}\,,$ (201)
while the contribution of internal vertices—the second product—is, by Proposition C.4,
$\prod_{v\in\tau:k_{v}(\tau)>0}R^{\frac{\sum_{l=1}^{k_{v}(\tau)}i_{vl}}{2}+1-\frac{j_{v}}{2}}P_{j_{v}}^{\left(\sum_{l=1}^{k_{v}(\tau)}(1-i_{vl}),\sum_{l=1}^{k_{v}(\tau)}i_{vl}\right)}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k_{v}(\tau)\right)\,.$ (202)
The exponents in $R$ have many cancellations and result in a global exponent of
$-k-i_{\varnothing}/2+\sum_{v\in\tau:k_{v}(\tau)=0}j_{v}/2=-k-\epsilon/2\,,$ (203)
since $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$. The rest of the contribution of $\tau$ is a polynomial of valuation $12\#\tau-$ 1 times $((R^{\prime}/R)^{2}-(S^{\prime}/\sqrt{R})^{2})^{-\#\tau}$. Multiplying and dividing by the determinant of $d_{\mathbf{t}}\mathbf{U}$ to the power $12k+$ 1, We finally obtain a contribution of $\tau$ of the form
$R^{-k-\epsilon/2}\frac{\Pi^{\tau}_{\pm,k}\left(\frac{R^{(l)}}{R},\frac{S^{(l)}}{\sqrt{R}},1\leq l\leq k\right)}{\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)^{2k+1}}\,,$ (204)
where $\Pi^{\tau}_{\pm,k}$ are polynomials of valuation $14k+$ 1. After multiplying by $R^{k}$, summing over all possible $\tau\in\mathcal{T}^{\mathbf{0}}_{\epsilon}(k)$, and applying (193) and the remark below the latter formula, we finally obtain that
$\bar{M}_{\pm,h}=R^{-1/2}\frac{\Pi_{\pm,h}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq h+1\right)}{\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)^{2h+1}}\,,$ (205)
[…]
$R^{h/2}M_{\pm,h}=\left(\frac{R^{\prime}}{R}\pm\frac{S^{\prime}}{\sqrt{R}}\right)\frac{\Pi_{\pm,h}\left(\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}},1\leq k\leq h+1\right)}{\left(\left(\frac{R^{\prime}}{R}\right)^{2}-\left(\frac{S^{\prime}}{\sqrt{R}}\right)^{2}\right)^{2h+1}}\,.$ (206)
Combining this with Theorem C.2, this concludes the proof that Proposition 4.8 holds for $g\geq 2$ and $n=0$.
Appendix D Higher order differentials of inverse functions of several variables
Here we provide a general discussion revisiting Lagrange inversion that is used in Section C.2.2. The results are probably known in some other form, the paper by Warren Johnson cited in [budd2020irreducible] pointing as far as Sylvester.
Fix $n\geq 1$ and let $\mathbf{x}=(x_{1},\ldots,x_{n})$, $\mathbf{y}=(y_{1},\ldots,y_{n})$ be formal variables. We let $\mathbf{f}(\mathbf{x})=(f_{1}(\mathbf{x}),\ldots,f_{n}(\mathbf{x}))$ and $\mathbf{g}(\mathbf{y})=(g_{1}(\mathbf{y}),\ldots,g_{n}(\mathbf{y}))$ be two families of $n$ elements in $K[[\mathbf{x}]]$ and $K[[\mathbf{y}]]$ respectively, where $K$ is some given field. We assume that $\mathbf{f}$ and $\mathbf{g}$ are compositional inverses, that is, $\mathbf{g}(\mathbf{f}(\mathbf{x}))=\mathbf{x}$ and $\mathbf{f}(\mathbf{g}(\mathbf{y}))=\mathbf{y}$. By classical consideration, given $\mathbf{f}$, the compositional inverse $\mathbf{g}$ exists if and only if $\mathbf{f}(\mathbf{0})=\mathbf{0}$ and the differential of $\mathbf{f}$ at $\mathbf{0}$ is invertible, which we assume.
Let us write the expansion
$\mathbf{f}(\mathbf{x})=a_{1}\mathbf{x}-a_{2}(\mathbf{x},\mathbf{x})-a_{3}(\mathbf{x},\mathbf{x},\mathbf{x})-\cdots\,,$ (207)
where $a_{1},-a_{2},-a_{3},\ldots$ are, up to a multiplicative coefficient, the successive differentials of $\mathbf{f}$ at 00, which are respectively symmetric $k$-linear, with the explicit expression $a_{1}(\mathbf{x})=\sum_{i=1}^{n}x_{i}\partial\mathbf{f}/\partial x_{i}(0)$, and
$-a_{k}(\mathbf{x},\ldots,\mathbf{x})=\frac{1}{k!}\sum_{i_{1},i_{2},\ldots,i_{k}\in\{1,\ldots,n\}}\frac{\partial^{k}\mathbf{f}}{\partial x_{i_{1}}\ldots\partial x_{i_{k}}}(0)x_{i_{1}}\cdots x_{i_{k}}\\ =\sum_{a_{1}+\cdots+a_{n}=k}\frac{\partial^{k}\mathbf{f}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(0)\frac{x_{1}^{a_{1}}}{a_{1}!}\cdots\frac{x_{n}^{a_{n}}}{a_{n}!}\,.$ (208)
This yields, after substituting $\mathbf{g}(\mathbf{y})$ to $\mathbf{x}$,
$\mathbf{g}(\mathbf{y})=a_{1}^{-1}\mathbf{y}+a_{1}^{-1}a_{2}(\mathbf{g}(\mathbf{y}),\mathbf{g}(\mathbf{y}))+a_{1}^{-1}a_{3}(\mathbf{g}(\mathbf{y}),\mathbf{g}(\mathbf{y}),\mathbf{g}(\mathbf{y}))+\cdots\,.$ (209)
This former expression has a natural tree interpretation, illustrated in Figure 7.
Figure 7: An illustration of the tree structure of derivatives of inverses.
We use the Ulam-Harris notation, where a tree $\tau$ is a subset of the set of integer words, rooted at the empty word $\varnothing$. Every vertex $v\in\tau$ has an arity (number of children) denoted by $k_{v}(\tau)$, and these children are the words $v1,v2,\ldots,vk_{v}(\tau)$. The trees considered are planted Pólya trees (with an extra edge pointing to the root vertex $\varnothing$), in which the arity $k_{v}(\tau)$ of every vertex $v\in\tau$ is an element of $\{0,2,3,4,\ldots\}$. We let $\mathcal{T}$ be the family of such trees in which the bottom and top half-edges of the parent edge of $v$ is decorated with two elements $i_{v},j_{v}\in\{1,\ldots,n\}$. For a given $\epsilon\in\{1,\ldots,n\}$, we also let $\mathcal{T}_{\epsilon}$ be the set of those trees for which $i_{\varnothing}=\epsilon$. Then
$g_{\epsilon}(\mathbf{y})=\sum_{\tau\in\mathcal{T}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v},j_{v}}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,,$ (210)
with the convention that $(a_{0})_{j}^{\varnothing}=y_{j}$, that is, the contributions of the variable $\mathbf{y}$ comes from the leaves of the trees. The partial derivatives of order $k$ of $g_{\epsilon}$ come from the finite family $\mathcal{T}_{\epsilon}(k)$ of elements of $\mathcal{T}_{\epsilon}$ with $k$ leaves. More precisely, for a given composition $\mathbf{a}=(a_{1},\ldots,a_{n})$ of $k$ with $n$ parts, one has
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(0)=\sum_{\tau\in\mathcal{T}^{\mathbf{a}}_{\epsilon}}\prod_{v\in\tau}(a_{1}^{-1})_{i_{v},j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}(a_{k_{v}(\tau)})_{j_{v}}^{i_{v1},i_{v2},\ldots,i_{vk_{v}(\tau)}}\,,$ (211)
where $\mathcal{T}^{\mathbf{a}}_{\epsilon}$ is the set of trees $\tau\in\mathcal{T}_{\epsilon}(k)$ with $a_{i}$ leaves $v$ labeled by $j_{v}=i$, for $1\leq i\leq n$.
By applying this reasoning to the series $\mathbf{f}(\mathbf{x}+\mathbf{x}^{\prime})-\mathbf{f}(\mathbf{x})$ and to $\mathbf{g}(\mathbf{y}+\mathbf{y}^{\prime})-\mathbf{g}(\mathbf{y})$ instead, we obtain
$\frac{\partial^{k}g_{\epsilon}}{\partial x_{1}^{a_{1}}\cdots\partial x_{n}^{a_{n}}}(\mathbf{f}(\mathbf{x}))=\sum_{\tau\in\mathcal{T}^{\mathbf{a}}_{\epsilon}(k)}\prod_{v\in\tau}(d_{\mathbf{x}}\mathbf{f}^{-1})_{i_{v},j_{v}}\prod_{v\in\tau:k_{v}(\tau)>0}\frac{-1}{k_{v}(\tau)!}\frac{\partial^{k_{v}(\tau)}f_{j_{v}}}{\partial x_{i_{v1}}\cdots\partial x_{i_{vk_{v}(\tau)}}}(\mathbf{x})\,.$ (212)
[…]